### v1

# Implementation Study of Cost-Effective Verification for Pietrzak’s Verifiable Delay Function in Ethereum Smart Contracts

Abstract
Verifiable Delay Function (VDF) is a cryptographic concept that ensures a minimum delay before output through sequential processing, which is resistant to parallel computing. Among the two well-known VDF protocols, Wesolowski and Pietrzak VDF, we focus on the Pietrzak VDF due to its computational efficiency and suitability for blockchain environments. Pietrzak’s approach uses a recursive proof verification with the halving protocol, offering a practical alternative despite the longer proof length than Wesolowski’s approach. Given the scarcity of research on practical VDF verification implementation, especially within smart contracts, this paper aims to implement cost-effective verification for the Pietrzak VDF in an Ethereum-based environment without compromising the VDF verification’s integrity and reliability. Firstly, we propose generalized proof generation and verification algorithms for potential efficiency improvement. Secondly, we categorize and measure the gas cost of each part in a transaction for VDF verification. Thirdly, based on the analysis, we theoretically predict the optimized proof construction. Finally, we demonstrate the theoretical prediction matches the implementation results. Furthermore, our research shows that the proof length of the Pietrzak VDF is generated under 8 KB with the security level of 2048 bits, much smaller than the previous expectation. This implies that the Pietrzak VDF can be practically used for cryptographic applications on blockchains.
§§footnotetext: These authors contributed equally to this work.
Keywords Ethereum  $\cdot$ blockchain security  $\cdot$ cryptographic protocol  $\cdot$ decentralized application  $\cdot$ gas optimization  $\cdot$ verifiable delay function
1 Introduction
footnotetext: We express our special thanks to Prof. Seungjoo Kim at Korea University for his invaluable advice on cryptographic algorithms.
In blockchain technology, one critical element is the generation and use of cryptographically secure random numbers. These secure random numbers play a pivotal role in various applications of blockchain; for example, in consensus, random numbers can ensure fair selections of sequencers and block proposers based on a probabilistic process. Random numbers are essential for ensuring anonymity and serve as the basis for cryptographic techniques. Despite the significance of secure random numbers, incorporating them into smart contracts results in challenges. Due to the lack of random number methods in the smart contract interface, individuals previously used to apply future block data (i.e. the blockhash of a future block) as an ad-hoc random number. It can be manipulated by a block generator. Hence, leveraging ad-hoc random numbers is not a suitable approach for the present situation.
In this context, two notable techniques have emerged: Verifiable Random Functions (VRFs) and Verifiable Delay Functions (VDFs). VRFs are cryptographic primitives that produce a random number along with proof of its correctness, allowing anyone to verify the number’s authenticity without compromising its unpredictability. Several commercial random number generating services using VRF, such as Chainlink VRF, are already available. As part of their operation, a node acting as an oracle transmits a randomly generated number from off-chain nodes to a blockchain.
On the other hand, VDFs are designed to produce a unique output that requires a predetermined amount of time to compute, but whose result can be efficiently verified by others. Therefore, VDFs do not directly generate a random number. Rather, VDFs support random number generation of the commit-reveal scheme by utilizing their time-lock property. The most critical weakness of the commit-reveal scheme is that all the participants must reveal to generate a secure random number. VDFs’ time-lock property ensures that the random outcome cannot be predicted or accelerated by computational power in case of any absence. It makes VDFs practical for applications that benefit from a guaranteed delay in randomness generation. Considering financial penalties do not perfectly guarantee the generation of random numbers, the commit-reveal scheme can enhance security with VDFs as a commit-reveal-recover scheme.
Figure 1: In our model, a user sends a VDF proof after the Pietrzak VDF evaluation. The primary issue is the significant gas usage for the VDF verification. The secondary issue is the length of the VDF proof as the Pietrzak VDF’s length increases as the time delay increases.
Pivotal advancements in VDFs have been made through the research of Pietrzak [1] and Wesolowski [2], who each proposed practical implementations of VDFs. Despite these theoretical developments, there remains a notable gap in research regarding the application of VDFs within blockchain environments. The challenge lies not only in the theoretical construction of VDFs, but also in adapting these constructions to the resource usage context of blockchain systems. For example, the Pietrzak VDF uses the halving protocol that makes the length of proof proportional to the delay time. It makes utilizing the Pietrzak VDF considered unrealistic in blockchain [3]. The Wesolowski VDF needs a hash-to-prime oracle that is not supported in most blockchains.
When implementing VDFs, the process can broadly be divided into three parts: evaluation, proof generation, and verification. In the context of blockchain, the verification component must be implemented on-chain. This segmentation is due to our belief that in typical blockchain applications, both evaluation and proof generation processes are performed off-chain, mainly due to their computational intensity and the associated costs when executed on a blockchain network.
Our goal has been to efficiently implement VDF verification within smart contracts as Figure 1’s illustration. Specifically, we have targeted the Pietrzak VDF model. There are two reasons: Firstly, the Wesolowski VDF requires a hash-to-prime function that is challenging to implement on smart contracts. Secondly, we conjectured that optimization can mitigate the long-proof length issue partially. Based on this algorithm choice, consequently, this research demonstrated that the Pietrzak VDF verification can be utilized in the EVM-compatible smart contract environment through various cost-saving analysis and experiments.
The contributions of this paper are as follows:
  • To the best of our knowledge, this research presents the first academic study of Pietrzak VDF implementation on smart contracts.
  • We proposed generalized algorithms for proof generation and verification of the Pietrzak VDF for potential efficiency enhancement.
  • Through theoretical analysis, we demonstrated the existence of a unique parameter that reduces the proof length. Additionally, the actual implementation results were consistent with our theoretical expectations.
  • Contrary to previous studies’ expectations, we demonstrated that the Pietrzak VDF could be utilized in the Ethereum environment with under 10 KB calldata size.
The rest of the paper is organized as follows: Section 2 introduces related studies about VDFs and their implementation. Section 3 describes the proposed target algorithms and our implementation environment. Section 4 illustrates the data structures of proofs and analyzes gas costs for the data and its dispatching cost. Section 5 studies arithmetic operations in the verification process and their gas costs. In Section 6, based on the previous sections’ outcomes, we present theoretical analysis and the cost-effective implementation. Section 7 discusses this research’s contributions, limitations, and future directions. Finally, Section 8 concludes the paper.
2 Background
This section introduces VDFs, cryptographic algorithms crucial for blockchain applications and secure timestamping. It also reviews related works that inform the VDF implementations discussed in this paper.
2.1 Verifiable Delay Functions
VDFs are cryptographic protocols crafted to ensure that their execution takes a predefined minimum amount of time, regardless of the computational power applied. Here we introduce the core mechanisms that define VDFs and the core properties that VDFs get from the elements.
2.1.1 Definition and Fundamental Properties
A VDF consists of three key algorithms that facilitate one round: setup, evaluation, and verification.
  • Setup $(\text{Setup}(1^{\lambda})\rightarrow(\mathbb{G},T))$: Initializes the VDF with public parameters. Here, $1^{\lambda}$ represents the security parameter indicating the required security level, $\mathbb{G}$ is a cryptographic group used in computations, and $T$ is the time delay parameter that sets the duration of the computation.
  • Evaluation $(\text{Evaluate}(\mathbb{G},T,X)\rightarrow(Y,\pi))$: For a given input $X$, the VDF computes an output $Y$ over time $T$, which is deliberately designed to be sequential, preventing acceleration through parallel processing. The output includes a proof $\pi$ that validates the computation.
  • Verification $(\text{Verify}(\mathbb{G},T,X,Y,\pi)\rightarrow\{\textsf{true},\textsf{false}\})$: Enables any observer to quickly and efficiently verify the correctness of the output $Y$ and its proof $\pi$ without undergoing the time-intensive computation themselves. A return of true confirms the validity of the output, while false indicates an error.
From the above algorithms, VDFs must get the three primary attributes [4]:
  • Sequentiality in evaluation: The computation must take a fixed minimum time, irrespective of the available computational resources.
  • Efficiency in verification: Although the computation itself is intentionally delayed, the verification process is designed to be quick and require minimal resources.
  • Uniqueness in output: Each input to a VDF yields a unique, unpredictable output, ensuring reliability and consistency in applications.
2.2 Feasibility of Applying VDFs to Smart Contracts
In both Wesolowski [2] and Pietrzak [1], the prover claims $y=x^{2^{T}}\mod N$ where N is a RSA modulus, but the way to verify it is different. In Wesolowski, the verifier gives the prover a random and large prime challenge value c, and the prover gives the verifier a proof $\pi$ which is $x^{{2^{T}}/c}$. Here, the prover does not compute $\pi$ anew but rather responds by taking the intermediate value of the process of computing $y=x^{2^{T}}\mod N$ from memory. The verifier then checks if $\pi\in\mathbb{Z}_{N}^{*}$ and accepts if $y={\pi^{c}}{x^{{2^{T}}\mod c}}$. To make this scheme non-interactive in a blockchain environment, the Fiat-Shamir heuristic is utilized: the prover computes $c$ directly as hash-to-prime(x, y) and passes it along with $\pi$ to the verifier in the first message. To ensure the difficulty in creating a verifiable proof, it is computationally assumed that calculating $z^{c}=y\mod N$ is challenging for any $z\neq 1$. Although this operation is time-independent with respect to $T$, it requires considerable resources and presents significant implementation challenges. These include incorporating an on-chain hash-to-prime function capable of handling relatively large data sizes, such as 2048 bits.
On the other hand, Pietrzak VDF does not require techniques like hash-to-prime and makes no computational assumptions. In the Pietrzak protocol, proof generation and verification are halving protocol that involves a loop $log(T)$ times, dividing $T$ by 2 until $T$ becomes 1. During each iteration of the loop in a proof generation, the prover computes ${v_{i}}=x^{2^{T/2}}$. The Fiat-Shamir heuristic is used to compute $r=Hash(x_{i},y_{i},v_{i})$ to make it non-interactive in a blockchain environment. And the prover and verifier calculate $(x_{i+1},y_{i+1})\leftarrow({{x_{i}}^{r}{v_{i}}},{{v_{i}}^{r}}{y_{i}})$ that satisfies $y_{x+i}=x_{i+1}^{2^{T/2}}$. Thus, a proof $\pi=\{v_{i}\}_{i\in[1,\log_{2}T]}$ of $\log_{2}T$ length is produced. When each element of the proof is over $2048$ bits, it is expected to be impractical for use in a blockchain due to the large size [3].
2.3 Related Works
In this subsection, we review the critical advancements in VDFs, from their theoretical foundations to practical implementations related to this study.
Theoretical research on VDFs. The concept of VDFs has garnered significant attention for its potential to enable new applications in cryptography and blockchain technology. The foundational work by Rivest, Shamir, and Wagner on "Time-Lock Puzzles and Timed-Release Crypto" [5] laid the groundwork for the development of protocols that secure data until a certain amount of time has passed. This seminal paper introduced the concept of time-lock puzzles, a precursor to VDFs, demonstrating how cryptographic operations could be used to create a delay mechanism that is practically infeasible to bypass without solving the puzzle.
Building upon these principles, Pietrzak [1] and Wesolowski [2] independently proposed two distinct approaches to VDFs, each with unique advantages in terms of efficiency and security. Pietrzak’s method focuses on using iterated sequential squaring in a group of unknown order, providing a balance between computation time and verification efficiency. Wesolowski’s approach, on the other hand, introduces an elegant solution that reduces the verification complexity, making it more practical for applications that require fast and efficient verification processes.
Further advancing the field, Dan Boneh et al. [6] have conducted extensive research on optimizing VDFs, addressing critical aspects such as the construction of more efficient protocols and the reduction of computational requirements for both generation and verification phases. Their work not only highlights the theoretical advancements in VDF technology but also emphasizes the practical implications for secure, decentralized systems.
Implementation of VDFs. The practical realization of VDFs has become an area of intense research and development, seeking to bridge the gap between theoretical robustness and real-world applicability. A significant contribution to this endeavor is the study by Attias et al. [3], which provides a comprehensive implementation analysis of two leading VDFs, Pietrzak and Wesolowski. Their work meticulously evaluates the performance of VDFs’ evaluation, proof generation, and verification. Their work provided us with insight into the research and optimization techniques for VDFs. However, they expected the Pietrzak VDF would not be viable for blockchains. This is contrary to our results that presented the feasibility of the Pietrzak VDF on Ethereum smart contracts.
Choi et al. [7] proposed three practical commit-reveal-recover techniques. Their protocols are crucial for establishing the commit-reveal-recover scheme, which mostly existed at the conceptual level. They incorporated the Wesolowski VDF into Ethereum smart contracts for the purpose of demonstration. Our research covers VDF verification, which corresponds to the Proof-of-Exponentiation (PoE) stated in their work. Their PoE gas usage is approximately 2.3 million, which is comparable to the gas consumption of our work with the time delay variable of $2^{22}$. Due to the Wesolowski VDF not having increased calculation requirements with greater time delays, our approach becomes less efficient when the time delay variable exceeds $2^{23}$. Nevertheless, the Wesolowski VDF encounters an issue when it comes to fully implementing the hash-to-prime function as previously indicated.
3 Implementation Target and Environment
In this section, we provide the target algorithm and environment. Firstly, we define the Pietrzak VDF algorithms we aim to implement in this research. Secondly, we describe the experiment environment in the following sections.
3.1 Pietrzak VDF Algorithms
Our research has a model with a smart contract that supports a function that performs Pietrzak VDF verification. Referring to the Pietrzak’s original paper [1], we define the evaluation as described in Algorithm 1 and the halving protocol for proof generation in Algorithm 2.
3.1.1 VDF Evaluation
$y\equiv x^{2^{T}}\equiv x^{2^{2^{\tau}}}\mod{N}$ (1)
The VDF evaluation process as defined in Algorithm 1 is pivotal in generating a VDF output from given inputs. It basically computes Equation 1. The algorithm takes an initial value $x$, a time parameter $T$ (also notated as $2^{\tau}$), and a modulus $N$ to produce the output $y$. $N$ that is a $\lambda$-bit number is produced by an RSA setup ($N=p\cdot q$ where $p,q$ are two big prime numbers). The process involves exponentiating $x$ to $2^{T}\mod N$ to ensure the difficulty and time consumption of the evaluation are as intended. This evaluation serves as the foundation for verifying that the output can only be computed with a predefined amount of real-time computational work, which is crucial in applications requiring guarantees against parallelization.
Algorithm 1 Evaluation of the Pietrzak VDF
1:input: $x,T,N$
2:output: $y$
3:$y\leftarrow x$
4:for $k\leftarrow 1$ to $T$ do
5:     $y\leftarrow y^{2}\mod N$
6:end for
7:return $y$
3.1.2 Halving Protocol
The halving protocol detailed in Algorithm 2 optimizes the proof generation process for VDFs by reducing the verification time incrementally. At each step, the protocol halves the time parameter $T$, recalculates intermediate values $x_{i}$ and $y_{i}$, and updates these values based on a cryptographic hash function to ensure integrity. As we implement it on Ethereum, we use Keccak256 as the cryptographic hash function. This iterative reduction continues until the desired granularity is reached, ensuring that the verifier’s workload is minimized while maintaining cryptographic security.
Algorithm 2 Halving Protocol of the Pietrzak VDF
1:input: $N,x_{i},T,y_{i}$
2:output: $N,x_{i+1},T/{2^{i}},y_{i+1},\mu_{i+1}$
3:$\mu_{i}\leftarrow x_{i}^{2^{T/{2^{i}}}}$
4:$r_{i}\leftarrow\text{Hash}(x_{i}+y_{i}+v_{i})$
5:$x_{i+1}\leftarrow x_{i}^{r_{i}}\cdot v_{i}\mod N$
6:if $T/2^{i-1}$ is odd then
7:     $y_{i}\leftarrow y_{i}^{2}\mod N$
8:end if
9:$y_{i+1}\leftarrow v_{i}^{r_{i}}\cdot y_{i}\mod N$
10:return $N,x_{i+1},T/{2^{i}},y_{i+1},\mu_{i+1}$
3.1.3 Introduction of Generalized Proof Generation/Verification and Its Rationale
In the original paper [1] and the related works [6, 3], the conventional Pietrzak halving protocol is repeated until the time parameter $T$ gets 1 for proof generation. It helps the verifier consume the nearly least resources to verify VDF proofs in general systems. However, on Ethereum, the cost of resource usage, called gas, significantly differs from general environments. Ethereum Yellowpaper [8] provides predefined gas prices for transactions and opcodes executions. Especially in this context, repeating the halving protocol till $T=1$ does not guarantee the most efficient verification. Therefore, we generalized the Pietrzak verification algorithm with flexible repetition of the halving protocol. In Algorithm 4, we introduced a new parameter $\delta$. The halving protocol is repeated till the time variable becomes equal to $\delta$. Then, the verifier checks $y\stackrel{{\scriptstyle?}}{{=}}x^{2^{2^{\delta}}}$.
3.1.4 Generalized Proof Generation
In Algorithm 3, the generalized proof generation is used to construct the cryptographic proofs required for VDF verification under flexible conditions. The proof generation starts with the evaluation result and halves the exponentiation of a proof step-by-step until reaching the specified parameter $\delta$. Each halving can be tracked as intermediate values $v_{i}$ are stored as part of the proof. This mechanism allows for flexible adjustment of the verification process to accommodate different system capacities and security requirements.
Algorithm 3 Generalized Proof Generation of the Pietrzak VDF with Evaluation
1:input: $x,T,\delta,N$
2:output: $\{\pi_{i}\}_{i=1}^{\tau-\delta}$
3:$\tau\leftarrow\lfloor\log_{2}(T)\rfloor$
4:$y\leftarrow x$
5:for $k\leftarrow 1$ to $T$ do
6:     $y\leftarrow y^{2}\mod N$
7:end for
8:$(x_{1},y_{1})\leftarrow(x,y)$
9:for $i\leftarrow 1$ to $\tau-\delta$ do
10:     $v_{i}\leftarrow x^{2^{T/2^{i}}}$
11:     $r_{i}\leftarrow keccak256(x_{i}+y_{i}+v_{i})$
12:     $x_{i+1}\leftarrow x_{i}^{r_{i}}\cdot v_{i}\mod N$
13:     if $T/2^{i-1}$ is odd then
14:         $y_{i}\leftarrow y_{i}^{2}\mod N$
15:     end if
16:     $y_{i+1}\leftarrow v_{i}^{r_{i}}\cdot y_{i}\mod N$
17:     $\pi_{i}\leftarrow v_{i}$
18:end for
19:return $\{\pi_{i}\}_{i=1}^{\tau-\delta}$
3.1.5 Generalized Verification
Algorithm 4 describes the generalized verification process for VDFs on Ethereum. It utilizes the proofs generated by the generalized proof generation method to validate the sequence of operations leading to the final result. The verifier iteratively computes and checks each step of the proof against the corresponding intermediate values and the final condition $y_{\tau-\delta+1}=x_{\tau-\delta+1}^{2^{2^{\delta}}}$. This process ensures that the computation was performed correctly and meets the specified delay constraints, which is crucial for applications relying on the integrity of timed cryptographic operations.
Algorithm 4 Generalized Verification of the Pietrzak VDF
1:input: $x,y,T,\delta,\{\pi_{i}\}_{i=1}^{\tau-\delta},N$
2:output: True or False
3:$\tau\leftarrow\lfloor\log_{2}(T)\rfloor$
4:$(x_{1},y_{1})\leftarrow(x,y)$
5:for $i\leftarrow 1$ to $\tau-\delta$ do
6:     $v_{i}\leftarrow\pi_{i}$
7:     $r_{i}\leftarrow keccak256(x_{i}+y_{i}+v_{i})$
8:     $x_{i+1}\leftarrow x_{i}^{r_{i}}\cdot v_{i}\mod N$
9:     if $T/2^{i-1}$ is odd then
10:         $y_{i}\leftarrow y_{i}^{2}\mod N$
11:     end if
12:     $y_{i+1}\leftarrow v_{i}^{r_{i}}\cdot y_{i}\mod N$
13:end for
14:if $y_{\tau-\delta+1}=x_{\tau-\delta+1}^{2^{2^{\delta}}}$ then
15:     return True
16:else
17:     return False
18:end if
According to Algorithm 3, and 4, the data should be submitted are the root ($x$), the VDF evaluation result ($y=x^{2^{T}}$), the time delay variable ($T$), the VDF proof ($\{\pi_{i}\}_{i=1}^{\tau-\delta}$), and the group description ($N$) as described in Equation 2.
$\Big{(}N,x,y,\tau,\overbrace{(\pi_{1},\pi_{2},\pi_{3},...,\pi_{k})}^{\overbrace{\lfloor\log_{2}(T)\rfloor}^{\tau}-\delta}\Big{)}$ (2)
3.2 Target Time Delay and the Bit Length
Recent advances in VDF implementation using C++, as studied by Attias et al. [3], demonstrate that VDF evaluation times can range from $2^{20}$ to $2^{25}$ cycles. These findings suggest that the implementations are not only effective but also efficient in experimental setups. In the context of Ethereum, where block generation typically requires around 12 seconds and even less time in other blockchain systems, we recommend that the time delay for generating VDF-based random numbers should not exceed one minute. This recommendation aims to prevent excessive resource consumption without compromising practical utility.
The experimental results from Attias et al. [3] show that a performance duration of approximately one minute at the 2048-bit level is feasible. Consequently, we advocate for a similar time range in VDF applications. The security of our target Pietrzak VDF configuration depends on the difficulty of the prime factoring problem, akin to that in the RSA encryption system. Currently, a 2048-bit key length is deemed sufficient for maintaining security standards. However, considering potential advancements in computational capabilities, a 3072-bit key length might be necessary after ten years. Therefore, our experiments included tests with both 2048-bit and 3072-bit settings for VDFs.
3.3 Environment for Gas Usage Measurement
Here we describe the setup of this research used to measure gas usage in Ethereum smart contracts, detailing the use of Solidity and the related framework.
3.3.1 Development Environment and Compiler Version
The development of the contract was conducted using Solidity version 0.8.23. This choice was driven by the compatibility and full support provided by the Hardhat framework, a comprehensive Ethereum development environment with integrated debugging capabilities. At the start of the development process, Solidity 0.8.23 was the latest compiler version fully supported by Hardhat for debugging and other essential development functionalities. Generally, gas optimization is best achieved with the latest version of the compiler.
3.3.2 Optimizer Configuration
For configuration, the core parameter of the Solidity optimizer is Runs which indicates how often each opcode will be used in the deployed contract. That is, higher Runs implies less gas cost for execution and higher gas cost for deployment. We set the optimizer parameter Runs to 4,294,967,295($2^{32}-1$), which is the maximum value of the parameter, indicating the highest level of optimization for execution.
We set up Via-IR (Intermediate Representation) and the Yul optimizer in the Solidity compiler to generate efficient bytecode. Via-IR helps efficiency, security, and simplicity in code generation by deviating from the direct compilation of Solidity code to EVM bytecode. Instead, via-IR processes an intermediate step where Yul code, an intermediate representation, is generated.
In this step, we can enhance the Yul optimizer’s efficiency. We replaced the msize() operation with the free memory pointer. This adjustment necessitated manual updates to the memory indexes, a process rigorously tested to ensure the integrity of our implementation. We also applied memory-safe assembly blocks in our code. This change not only aligns with Solidity’s memory model but also activates Via-IR in the compiler settings, resulting in significant gas savings.
3.3.3 Gas Measurement
We implemented external functions in MinimalApplication.sol, the major source file for implementation, to measure the gas used with different strategies. The gas cost depends on the order of the function selectors in these external functions, but the differences are considered to be negligible, with a maximum difference of about 60 gas used across the seven external functions. Therefore, an error margin of about 60 gas exists for the results presented in the figures.
3.3.4 Local Testing Environment
We utilized the Hardhat node, provided by the Hardhat framework, to create a local testing environment. The amount of gas usage was obtained using the eth_estimateGas method, which generates and returns an estimate of the gas required to complete the transaction. This method accurately measures the gas used in the execution of different contract functions and provides valuable insights into the efficiency of the implementations. It is important to note that the actual gas cost incurred on the real network is determined by multiplying the estimated gasUsed by the gasPrice at the Ethereum network circumstance.
4 Data and Accompanying Gas Cost
In this section, we introduce the data structure we used to send VDF proofs in the Ethereum transaction calldata. Then, we investigate the gas cost to send VDF proofs and regress the results to linear functions for the next sections.
4.1 Data Structure
In this study, the VDFs are implemented with key sizes of 2048 and 3072 bits. However, the EVM natively supports arithmetic operators only up to 256-bit numbers. Therefore, to facilitate efficient computation of numbers exceeding 256 bits, it is necessary to employ customized data formats. We revised the big number library implementation for Solidity developed by Firo [9]. The original data structure in this library includes bytes data (bytes val), bit length (uint bitlen), and a sign indicator (bool neg). Among the three elements, we deleted the sign indicator and its related low-level logic since our implementation does not require handling negative values. This library optimizes on-chain operations such as addition and subtraction by adding leading zeros for 32-byte alignment with the EVM’s memory word size, eliminating offset management. Also, it utilizes the bit length data field to optimize multiplication and comparison operations. Therefore, we format the bytes data and pre-calculate bit length off-chain before submission, significantly reducing unnecessary on-chain computations. For example, the data format for a 2048-bit representation in JSON is as shown in Listing 1. In this way, we can reduce the gas cost even though the calldata size increases. This decision is reasonable because the primary goal is to reduce the gas cost, not the transaction calldata size.
1{
2 "big_number": {
3 "val": "0x4621c26320fe0924bba1b7d5bc863495b9f0db3823b12a9a18e21d23...........b2127fd6390f5f6234a164bd39d1dd6884b768c4dd790586ee",
4 "bitlen": 2047
5 }
6}
Listing 1 Calldata for a Big Number Example
4.2 Gas Cost of Calldata
The gas cost for Ethereum transaction calldata was fixed in EIP-2028 "Transaction data gas cost reduction" [10]. According to the document, the cost of calldata is divided into two cases:
  • Non-zero byte: 16 gas
  • Zero byte: 4 gas
Considering the above gas pricing, we can calculate the cost for sending one 2048-bit length data in the aforementioned data structure. 2048-bit data is transformed into 512-length hex letters.
In our implementation, the function verifyRecursiveHalvingProof processes the generalized verification (Algorithm 4). It is defined as the Listing 2 in Solidity codes.
1function verifyRecursiveHalvingProof(
2 BigNumber[] memory v,
3 BigNumber memory x,
4 BigNumber memory y,
5 BigNumber memory n,
6 uint256 delta,
7 uint256 T
8)
Listing 2 Interface of the Halving Verification Function
A call to this function involves encoding parameter values. Start by computing the keccak256 of the UTF-8 byte representation of the string111verifyRecursiveHalvingProof((bytes,uint256)[],(bytes,uint256),(bytes,uint256),(bytes,uint256),uint256,uint256) based on Listing 2 and use the first four bytes, which results in 0xd8e6ac60.
For static types such as uint256, fields including delta and T are padded to 32 bytes. For dynamic types including bytes, struct, and dynamic arrays, they are encoded by specifying an offset to the start of its data area, followed by the length of the data and the actual data. In case of BigNumber[] memory v, it is the dynamic array of dynamic type struct BigNumber that contains dynamic bytes val. So the way dynamic types are encoded is recursively applied to each halving proof v.
In EVM, dynamic types such as arrays and structs are encoded by specifying an offset to the start of the data, its length, and then the data itself. This method applies recursively for nested dynamic types. Parameters such as $T$, $\delta$, $\lambda$, and parameter count (e.g., 6 in our model) determine the offsets and lengths. The gas usage for data depends probabilistically on the distribution of zero and non-zero bytes in the data. Since EVM organizes memory in 32-byte words, we can consider the same number of word is repeated when data is repeated. Finally, we can predict the approximate gas usage of calldata using this EVM data construction structure.
Considering the precision of data and its cost, we performed experiments to test the size of calldata and the related gas usage. The related gas usage comprises of the intrinsic gas and data dispatching gas. The intrinsic gas in EVM refers the minimum cost for a transaction. It contains the transaction cost and the cost for data allocation. For verification, we need not only to allocate data in a transaction, but also to dispatch data into the verification function. The dispatching cost is proportional to the data size. Therefore, we decided to include the dispatching cost into the data cost in this study. Figure 2(a) for $\lambda=2048$ and 2(b) for $\lambda=3072$ show the calldata size with the number of proofs and the related gas usage. We can see the gas usage and the calldata size simply and linearly increase with the number of proofs.
Figure 2: Experiment results of the calldata length and cost for the proofs
(a) Calldata size and gas cost including intrinsic gas and dispatching gas ($\lambda=2048$)
(b) Calldata size and gas cost including intrinsic gas and dispatching gas ($\lambda=3072$)
Finally, to elucidate the relationship between transaction size and gas costs, we perform a regression analysis. By applying the least squares method to approximate the best fit, we derive the following linear regression equations 3 and 4 for data sizes of 2048 and 3072 bytes, respectively:
$\mathcal{G}_{data,2048}(\tau)=5295.73\cdot\tau+37610.11$ (3)
$\mathcal{G}_{data,3072}(\tau)=7389.31\cdot\tau+43943.88$ (4)
These equations estimate the gas cost $\mathcal{G}$ as a function of transaction size $\tau$, highlighting the increase in required gas with larger transaction calldata.
5 Arithmetic Operations and Accompanying Gas Cost
This section focuses on analyzing the expenses associated with performing arithmetic operations in the context of verifying the Pietrzak VDF. The cost consists of two components: the cost of repeated halving and the cost of modular exponentiation for the final check. The algorithm executes a halving operation and then proceeds with a modular exponentiation verification. The modular exponentiation cost is studied before the halving cost, as the halving process includes the modular exponentiation. Using the results, we perform a regression analysis to formulate mathematical equations for the cost in the section that follows.
5.1 Modular Exponentiation Cost
This part explores strategies for optimizing the computational cost of modular exponentiation, a frequent operation in various cryptographic algorithms. It discusses algorithms such as Square and Multiply [11], Dimitrov multiexponentiation [12], and the ModExp contract introduced by EIP-198.
5.1.1 Square and Multiply
The square and multiply algorithm, also called binary exponentiation, is an efficient approach for computing exponentiation operations $x^{n}$, particularly beneficial for large number modular exponentiation. As described in Algorithm 5, by converting the exponent $b$ into its binary representation, the algorithm optimizes computation through a series of square (squaring) and multiply (multiplying by the base $x$) operations based on the binary digits of $n$. For an exponent represented by $k$ bits, the algorithm requires at most $k-1$ squaring and $k-1$ multiplication operations, making the computational complexity $O(\log_{2}n)$. This optimization is simple but significantly enhances efficiency for large-scale modular exponentiation tasks.
Algorithm 5 Square and Multiply for Modular Exponentiation
1:input: $a,b,N$
2:output: $a^{b}$ mod $N$
3:$r\leftarrow 1$
4:$a\leftarrow a\mod N$
5:while $b>0$ do
6:     if $b\mod 2=1$ then
7:         $r\leftarrow(r\cdot a)\mod N$
8:     end if
9:     $b\leftarrow b>>1$
10:     $a\leftarrow(a\cdot a)\mod N$
11:end while
12:return $r$
5.1.2 Dimitrov Multiexponentiation
The Dimitrov algorithm [12, 3] presented in Algorithm 6 enhances the efficiency of large integer multiplication, crucial in fields such as cryptography and computational mathematics, by significantly reducing the computational complexity. Traditionally, the multiplication of large integers ($x^{a}y^{b}$) exhibits a complexity of $O(2\log_{2}max(a,b))$. Utilizing the binary exponentiation method [13], the Dimitrov algorithm achieves a computational complexity closer to $O(\frac{7}{4}\log_{2}max(a,b))$. This characteristic allows for the rapid and resource-efficient multiplication of very large integers, making the algorithm particularly valuable for applications requiring efficient large-scale numerical operations. In the generalized verification (Algorithm 4), lines 9 – 12 perform multiexponentiation only if $T/2$ is odd. Therefore, we considered the Dimitrov algorithm as a candidate algorithm to apply for optimization.
Algorithm 6 Dimitrov Multiexponentiation Algorithm
1:input: $x,y,a,b,N$
2:output: $x^{a}\cdot y^{b}\mod N$
3:$h\leftarrow\max(\lfloor\log_{2}a\rfloor+1,\lfloor\log_{2}b\rfloor+1)$
4:$z\leftarrow 1$
5:$q\leftarrow x\cdot y\mod N$
6:for $i=h-1$ down to 00 do
7:     $z\leftarrow z\cdot z\mod N$
8:     if $a_{i}=1$ and $b_{i}=0$ then
9:         $z\leftarrow z\cdot x\mod N$
10:     else if $a_{i}=0$ and $b_{i}=1$ then
11:         $z\leftarrow z\cdot y\mod N$
12:     else if $a_{i}=1$ and $b_{i}=1$ then
13:         $z\leftarrow z\cdot q\mod N$
14:     end if
15:end for
16:return $z$
Figure 3: Experiment results of the modular exponentiation cost
(a) Gas usage of exponentiation by squaring, square and multiply, and precompiled ModExp, x is 2048-bit level
(b) Gas usage of exponentiation by squaring, square and multiply, and precompiled ModExp, x is 3072-bit level
(c) Gas usage of Dimitrov multiexponentiation and only precompiled ModExp, x is 2048-bit level
(d) Gas usage of Dimitrov multiexponentiation and only precompiled ModExp, x is 3072-bit level
5.1.3 ModExp by EIP-198
EIP-198 [14], known for introducing "Big Integer Modular Exponentiation", introduces the ModExp (Modular Exponential) precompiled contract to the EVM, offering a streamlined and cost-efficient approach to conducting large integer modular exponentiation, a foundational operation in numerous cryptographic algorithms.
This function calculates $a^{b}\mod n$, where $a$, $b$, and $n$ are large numbers, and prior to EIP-198, such execution on-chain was not officially supported by any operations. The ModExp contract mitigates these expenses by providing a standardized computation method, thus facilitating the incorporation of cryptographic computations within smart contracts. Building on this, EIP-2565 [15] contributed to further refining the gas cost model for these modular exponentiation operations by proposing adjustments to the pricing formula.
5.1.4 Comparison of Exponentiation Techniques and Regression
Figure 3(a) and Figure 3(b) present the gas usage experiment results of three modular exponentiation techniques. We name a naive exponentiation approach as exponentiation by squaring in the figures. Even if square and multiply (Algorithm 5 reduces the gas usage, it is obvious that the precompiled ModExp uses outstandingly less gas for the modular exponentiation than the others. Figure 3(c) and Figure 3(d) show the gas usage experiment results for modular multiexponentiation. We implemented Dimitrov multiexponentitaion including the precompiled ModExp. The comparing data named ‘Precompile’ is the gas usage when only the precompiled ModExp is used for multiexponentiation. The latter takes significantly less gas than the Dimitrov algorithm. Therefore, we applied ModExp for both the modular exponentiation and multiexponentiation operations in the implementation in Algorithm 4.
In the experiment, the pricing of ModExp is linearly proportional to the number of exponentiation as $\mathcal{G}(T)=aT+b$. As Equation 1, we denote $\tau$ as the logarithmic value of T. We could get the following results from a simple regression process.
$\mathcal{G}_{exp,2048}(\tau)=346.92\cdot 2^{\tau}+33432.08$ (5)
$\mathcal{G}_{exp,3072}(\tau)=780.83\cdot 2^{\tau}+37445.60$ (6)
5.2 Halving Cost
The halving cost of the arithmetic operations in the verification indicates the gas usage for lines 5 – 13 in Algorithm 4. The modular exponentiation operations in lines 8, 10, and 12 use the precompiled ModExp since its best efficiency in the previous subsection. For simplicity, as we assumed $T=2^{\tau}$, in this case, the if statement in line 9 is always false, therefore line 10 is never executed. Hence, the gas usage will be clearly proportional to the number of the repetition time, that is, $\tau-\delta$.
Figure 4: Experiment results of the halving repetition cost
(a) Gas usage of the halving algorithm ($\lambda=2048$)
(b) Gas usage of the halving algorithm ($\lambda=3072$)
The results of our experiments demonstrate the anticipated correlation. Figure 4(a) and Figure 4(b) display the gas cost of the repeated halving operations with $\lambda=2048$ and $\lambda=3072$ correspondingly. The data exhibits a clear linear relationship, allowing us to perform a regression analysis and model the results using a linear function. By employing the method of least squares approximation, we obtain the subsequent functions as Equation 7 and 8:
$\mathcal{G}_{halving,2048}(\tau)=165482.26\cdot\tau-155642.3$ (7)
$\mathcal{G}_{halving,3072}(x)=306089.32\cdot\tau-238572.65$ (8)
6 Cost-Effective Implementation
This section explains the implementation of a cost-efficient Pietrzak VDF verifier as an Ethereum smart contract. In the beginning, we utilize the gas cost outcomes obtained from the preceding sections to make theoretical predictions for the optimized proof generating process. Secondly, we prove that there exists a unique $\delta$ that minimizes the gas usage for VDF verification, and this minimization is independent of $\tau$. In the end, we compare the anticipated outcome based on theory with the actual results obtained from the experiment.
6.1 Theoretical Analysis on Proof Generation for Gas Usage Optimization
The optimized proof generation in this research means minimizing the verification gas cost. We can define the total gas cost function $g_{total}$ using the previous regression results.
$\mathcal{G}_{total}(\tau,\delta)=\mathcal{G}_{calldata}+\mathcal{G}_{halving}+\mathcal{G}_{exp}+C$ (9)
$=\big{(}\alpha\cdot(\tau-\delta)+c_{1}\big{)}+\big{(}\beta\cdot(\tau-\delta)+c_{2}\big{)}+\big{(}\gamma\cdot 2^{\delta}+c_{3}\big{)}+C$ (10)
$=(\alpha+\beta)(\tau-\delta)+\gamma\cdot 2^{\delta}+C^{\prime}$ (11)
Then, we get Theorem 1 says that there is a unique delta minimizing the total gas cost.
Theorem 1.
The function $\mathcal{G}_{total}(\tau,\delta)$ has exactly one unique minimum for $\delta$ when $\tau>\delta=\log_{2}(\alpha+\beta)-\log_{2}(\ln{2}\cdot\gamma)$.
Proof.
Firstly, we calculate the partial derivative of $\mathcal{G}_{total}$ with respect to $\delta$, which yields
$\frac{\partial\mathcal{G}_{total}}{\partial\delta}=\ln{2}\cdot\gamma\cdot 2^{\delta}-(\alpha+\beta).$
To show that there is a unique $\delta$ that minimizes $\mathcal{G}_{total}$, we need to find where this derivative equals zero. This leads us to the equation
$2^{\delta}=\frac{\alpha+\beta}{\ln{2}\cdot\gamma}.$
The function $2^{\delta}$ is strictly increasing. Hence, the equation $2^{\delta}=\frac{\alpha+\beta}{\ln{2}\cdot\gamma}$ has a unique solution $\delta$, which can be expressed as
$\delta=\log_{2}(\alpha+\beta)-\log_{2}(\ln{2}\cdot\gamma).$
This $\delta$ is where the first derivative equals zero.
Further, to confirm that this $\delta$ is indeed a minimum, we need to check the second derivative. The monotonic nature of the exponential function in this context implies that as $\delta$ increases, $\gamma\cdot 2^{\delta}$ grows without bounds, while decreasing $\delta$ below this point leads to $\gamma\cdot 2^{\delta}$ becoming smaller than $\alpha+\beta$, thereby confirming that the function is minimized at this $\delta$.
Thus, $\mathcal{G}_{total}(\tau,\delta)$ indeed has one unique minimum at $\delta$ within the specified range of $\tau>\delta\geq 0$, completing the proof. ∎
Also, we denote the minimizing $\delta$ as
$\delta_{m}=\arg\min_{\delta}\mathcal{G}_{total}(\tau,\delta).$ (12)
From Theorem 1, we can get the unique $\delta_{m}$ optimizing the gas cost. Applying all the regression results in Equation 3, 4, 5, 6, 7, and 8, we get the optimizing $\delta_{m}\approx 9.47$ for $\lambda=2048$ and $\delta_{m}\approx 9.18$ for $\lambda=3072$ respectively. In practice, $\delta$ for implementation must be an integer. Therefore, we can get the practical integer $\delta$ with the following corollary.
Corollary 2.
Let $\delta_{m}$ be a real number, and let $\mathcal{G}(\tau,\delta)$ denote a function in $\tau$ and $\delta$. Define $\Delta=\delta_{m}-\lfloor\delta_{m}\rfloor$ as the fractional part of $\delta_{m}$. Then, the relationship between $\mathcal{G}(\tau,\lfloor\delta_{m}\rfloor)$ and $\mathcal{G}(\tau,\lceil\delta_{m}\rceil)$ depends on $\Delta$ as follows:
  1. 1.
    If $\Delta<-\log_{2}(\ln 2)$, then $\mathcal{G}(\tau,\lfloor\delta_{m}\rfloor)<\mathcal{G}(\tau,\lceil\delta_{m}\rceil)$.
  2. 2.
    If $\Delta=-\log_{2}(\ln 2)$, then $\mathcal{G}(\tau,\lfloor\delta_{m}\rfloor)=\mathcal{G}(\tau,\lceil\delta_{m}\rceil)$.
  3. 3.
    If $\Delta>-\log_{2}(\ln 2)$, then $\mathcal{G}(\tau,\lfloor\delta_{m}\rfloor)>\mathcal{G}(\tau,\lceil\delta_{m}\rceil)$.
Proof.
Define
$f(\delta)=\mathcal{G}(\tau,\delta+1)-\mathcal{G}(\tau,\delta)=\gamma\cdot(2^{\delta+1}-2^{\delta})+(\alpha+\beta)(-(\delta+1)+\delta)=\gamma\cdot 2^{\delta}-(\alpha+\beta).$ (13)
Then, $f(\delta)=0$ when
$\delta=\log_{2}(\alpha+\beta)-\log_{2}\gamma.$ (14)
Also, as given $\gamma>0,\delta>0$
$\frac{df}{d\delta}=\ln 2\cdot\gamma\cdot 2^{\delta},$ (15)
$f$ is a monotone-increasing function in the boundary. Therefore, after the point $\delta^{\prime}$ where
$\delta^{\prime}=\log_{2}(\alpha+\beta)-\log_{2}\gamma=\delta_{m}+\log_{2}(\ln 2),$ (16)
$f$ is bigger than 0.
For the 1st case of the lemma, $\Delta<-\log_{2}(\ln 2)$ implies
$f(\lfloor\delta_{m}\rfloor)>0.$ (17)
As the definition of $f$,
$\mathcal{G}(\tau,\lfloor\delta_{m}\rfloor)<\mathcal{G}(\tau,\lceil\delta_{m}\rceil).$ (18)
In the same way, we can get the other two cases. ∎
Since $\log_{2}(\ln 2)\approx-0.52$, the optimal integer value of $\delta$ is 9 for both $\lambda=2048$ and $3072$, as the 1st case of Corollary 2.
6.2 Optimized Verification Gas Cost
Figures 5(a) and 5(b) depict the total gas cost for Pietrzak VDF verification at $\lambda=2048$ and $\lambda=3072$ respectively. Consistent with the theoretical predictions in Theorem 1, the gas usage graph shows a decline to a minimal point before increasing sharply. Optimal gas costs were observed at $\delta=9$ for both configurations, aligning perfectly with the findings of Corollary 2. This alignment between theoretical expectations and empirical results confirms the accuracy of our model and highlights its potential for optimizing Pietrzak VDF verification in blockchain applications.
Table 1 details the gas consumption and calldata sizes for the implemented parameters, uniformly using $\delta=9$. Contrary to the previous expectation of calldata sizes around 40KB [3], the sizes recorded were significantly lower, ranging from 5.52 to 7.35 KB for $\lambda=2048$ and from 7.22 to 9.72 KB for $\lambda=3072$. Gas usage varied from 1.96 to 2.80 million for $\lambda=2048$ and from 3.68 to 5.26 million for $\lambda=3072$. These findings not only demonstrate the efficacy of our approach but also highlight potential areas for further optimization in VDF implementations. Such improvements could lead to significant cost reductions in blockchain networks, enhancing the overall efficiency of cryptographic operations.
Figure 5: Experiment results for halving proof skipping parameter $\delta$ with $\tau=22,23,24$
(a) Skipping parts of halving proofs ($\lambda=2048$)
(b) Skipping parts of halving proofs ($\lambda=3072$)
Table 1: Implementation results based on the analysis
$\lambda$ $\delta$ $\tau$ Gas used Calldata size (KB)
2048 9 20 1970844 5.38
21 2133372 5.75
22 2297629 6.13
23 2470594 6.50
24 2620907 6.89
25 2799805 7.25
3072 9 20 3680363 7.13
21 3967114 7.63
22 4318050 8.13
23 4614955 8.63
24 4947040 9.13
25 5254521 9.63
7 Discussion
This section discusses the implementation of VDFs on Ethereum, focusing on programming choices for Ethereum’s Virtual Machine and their impact on gas optimization. We introduce challenges associated with VDF deployment on Ethereum’s Layer 1 due to high costs and explore Layer 2 solutions. The discussion concludes with an analysis of economic factors critical for VDF integration and sustainability on the blockchain.
7.1 Implementation Languages
There are main development languages in EVM-compatible network environments, for example, Viper, Huff, Solidity, and Yul. For our implementation, which requires dynamically adjusting proof lengths, we excluded Vyper due to its inability to handle dynamic data structures efficiently. Instead, we chose Solidity combined with Yul, leveraging Solidity for its comprehensive features and developer familiarity, and Yul for its superior gas optimization capabilities. This combination provides a balance between high-level language safety and low-level control over the Ethereum Virtual Machine.
Considering future enhancements, we are exploring Huff, noted for its exceptional gas efficiency and direct control over the EVM. Huff originated from the Aztec team where contracts needed extreme efficiency, even more than what Yul could provide. Although Huff offers significant optimization opportunities, it lacks the built-in safety checks of higher-level languages, raising the implementation complexity. Nevertheless, given our implementation’s emphasis on gas optimization, the benefits of adopting Huff justify its consideration for future development.
7.2 Practicality of VDF on EVM
Implementing VDFs on Ethereum’s Layer 1 presents significant challenges primarily due to high gas costs and the computational intensity required. Our detailed optimization trials reveal that despite employing cutting-edge optimization techniques, verifying a Pietrzak VDF on Layer 1 still demands at least one million gas units. This high resource consumption makes VDF applications impractical for widespread use on Ethereum’s primary network, especially for operations requiring frequent and repetitive VDF verifications. The economic implications of these requirements can stifle innovation and discourage developers from incorporating VDFs into their decentralized applications, limiting the exploration of potentially valuable use cases such as randomized governance protocols or secure public randomness in blockchain networks.
However, the emergence of Layer 2 solutions, such as optimistic rollups and zkRollups, offers a promising alternative for leveraging VDFs within the Ethereum ecosystem. These technologies provide a secondary processing layer that can handle complex computations off the main Ethereum chain, thereby significantly reducing gas costs and improving transaction throughput. Platforms like Optimism capitalize on these advantages, offering gas prices that are often thousands of times cheaper than those on Layer 1. This dramatic cost reduction opens up new possibilities for using VDFs on Ethereum, allowing for more frequent and economically viable implementations. Layer 2 solutions not only mitigate the high cost barriers associated with Layer 1 but also broaden the scope for adopting VDFs in a variety of applications, thus encouraging further development and research into this promising area of blockchain technology.
7.3 Lack of Economic VDF Protocols in Ethereum
The integration of VDFs into smart contracts holds great promise for enhancing blockchain security and reliability. However, beyond the technical challenges, there exists a substantial hurdle in the lack of a robust economic model specifically designed to support the unique operational needs of VDFs, such as incentivizing both provers and verifiers and managing commit-reveal-recover protocols. The absence of such an economic framework complicates the practical deployment of VDFs as it does not adequately address compensation for participants who dedicate substantial computational resources to generate and verify proofs. This gap can deter the participation of potential contributors, ultimately hindering the scalability and effectiveness of VDFs in real-world blockchain applications.
Moreover, considering the decentralized essence of blockchain technology, it is crucial that any protocol, especially those as intricate as VDFs, is underpinned by an economic strategy that not only supports technical execution but also ensures economic sustainability without the need for central governance. This calls for innovative research into economic models that are specifically tailored to the decentralized context of blockchain networks. These models need to effectively balance incentives and deterrence, fostering a system where honest participation is rewarded while dishonest behavior is penalized. Such economic strategies are vital for maintaining the integrity and fairness of the blockchain ecosystem, ensuring that VDFs are not only technically feasible but also economically viable and beneficial to all participants, thereby aligning with the overarching principles of decentralization and trust minimization in blockchain systems.
7.4 Future Works on a General Environment
Through this research, we investigated the efficient implementation of Pietrzak VDF verification within the Ethereum Virtual Machine (EVM) environment and observed results that align both theoretically and experimentally. During our study, we identified significant effects of generalized proof generation and verification, which were also included within the predictable realms of theoretical analysis. We anticipate that such an approach could be applicable not only within the EVM but also in more general computing environments. Although resource pricing in general systems would differ from the gas costs in the EVM, with elements like the repetition of halving protocols and resource consumption proportional to the proof length being predictable, the costs of exponentiation in VDF ($2^{2^{T}}$) are expected to be exponential. Therefore, based on the outcomes of this study, further research on optimizing Pietrzak VDFs in general computing environments appears reasonable.
8 Conclusion
In conclusion, our research has successfully demonstrated the feasibility and significance of implementing VDFs within blockchain environments, particularly through smart contracts in Ethereum. By focusing on the Pietrzak VDF construction, we have not only bridged a crucial gap in the academic studies of VDFs but also highlighted their practical utility in enhancing the security and efficiency of blockchain applications. Contrary to previous expectations, our work represents the first academic result in the cost-effective on-chain verification of the Pietrzak VDF, showing significant potential in ensuring cryptographically secure random number generation across various blockchain applications. Through strategic implementation, we have achieved substantial cost reductions, making the Pietrzak VDF more accessible and applicable. This research opens new avenues for the integration of VDFs in blockchain systems, promising to enhance security, fairness, and functionality in the evolving distributed ledger technologies.

### v2

@@
   • Uniqueness in output: Each input to a VDF yields a unique, unpredictable output, ensuring reliability and consistency in applications.
-2.2 Feasibility of Applying VDFs to Smart Contracts
-In both Wesolowski [2] and Pietrzak [1], the prover claims $y=x^{2^{T}}\mod N$ where N is a RSA modulus, but the way to verify it is different.
-In Wesolowski, the verifier gives the prover a random and large prime challenge value c, and the prover gives the verifier a proof $\pi$ which is $x^{{2^{T}}/c}$.
-Here, the prover does not compute $\pi$ anew but rather responds by taking the intermediate value of the process of computing $y=x^{2^{T}}\mod N$ from memory.
-The verifier then checks if $\pi\in\mathbb{Z}_{N}^{*}$ and accepts if $y={\pi^{c}}{x^{{2^{T}}\mod c}}$.
-To make this scheme non-interactive in a blockchain environment, the Fiat-Shamir heuristic is utilized: the prover computes $c$ directly as hash-to-prime(x, y) and passes it along with $\pi$ to the verifier in the first message.
-To ensure the difficulty in creating a verifiable proof, it is computationally assumed that calculating $z^{c}=y\mod N$ is challenging for any $z\neq 1$.
-Although this operation is time-independent with respect to $T$, it requires considerable resources and presents significant implementation challenges.
-These include incorporating an on-chain hash-to-prime function capable of handling relatively large data sizes, such as 2048 bits.
-On the other hand, Pietrzak VDF does not require techniques like hash-to-prime and makes no computational assumptions.
-In the Pietrzak protocol, proof generation and verification are halving protocol that involves a loop $log(T)$ times, dividing $T$ by 2 until $T$ becomes 1.
-During each iteration of the loop in a proof generation, the prover computes ${v_{i}}=x^{2^{T/2}}$.
-The Fiat-Shamir heuristic is used to compute $r=Hash(x_{i},y_{i},v_{i})$ to make it non-interactive in a blockchain environment.
-And the prover and verifier calculate $(x_{i+1},y_{i+1})\leftarrow({{x_{i}}^{r}{v_{i}}},{{v_{i}}^{r}}{y_{i}})$ that satisfies $y_{x+i}=x_{i+1}^{2^{T/2}}$.
-Thus, a proof $\pi=\{v_{i}\}_{i\in[1,\log_{2}T]}$ of $\log_{2}T$ length is produced.
-When each element of the proof is over $2048$ bits, it is expected to be impractical for use in a blockchain due to the large size [3].
+2.2 Feasibility of On-Chain VDF Verification
+Both Wesolowski [2] and Pietrzak [1] VDFs operate by transforming a value $x$ into $y=x^{2^{T}}\mod N$, where $N$ is an RSA modulus, although their verification methods differ.
+Scalability and implementation costs are significant challenges in blockchain systems.
+Employing the Fiat-Shamir heuristic, both methods facilitate non-interactive verification, suitable for blockchain environments.
+However, the computational intensity and storage demands for managing large numeric values pose substantial hurdles.
+Wesolowski’s VDF maintains a constant proof length ($O(1)$), independent of the time delay parameter $T$, making it particularly advantageous where proof size is a concern.
+The implementation on blockchain platforms like Ethereum is particularly challenging due to the high cost associated with verifying or finding primes for large numbers, such as 2048 bits.
+This necessity significantly increases the implementation cost, making full on-chain execution unfeasible.
+Pietrzak’s VDF does not require operations like hash-to-prime, simplifying its on-chain implementation.
+The protocol involves a halving process, where $T$ is divided by two until it becomes one, reducing the complexity of computations with each iteration.
+However, the proof length and cost are influenced by the time delay variable $T$, and the proof length grows with $O(\log_{2}(T))$.
+This characteristic can result in verifications that exceed the size limit for feasible blockchain implementation.
+The research anticipated that the length of the proof will approximate 40 KB [3].
 2.3 Related Works

### v3

@@
 Finally, we demonstrate the theoretical prediction matches the implementation results.
-Furthermore, our research shows that the proof length of the Pietrzak VDF is generated under 8 KB with the security level of 2048 bits, much smaller than the previous expectation.
+Furthermore, our research shows that the proof length of the Pietrzak VDF is generated under 8 KB with the 2048-bit RSA key length, much smaller than the previous expectation.
 This implies that the Pietrzak VDF can be practically used for cryptographic applications on blockchains.
@@
 It makes VDFs practical for applications that benefit from a guaranteed delay in randomness generation.
-Considering financial penalties do not perfectly guarantee the generation of random numbers, the commit-reveal scheme can enhance security with VDFs as a commit-reveal-recover scheme.
+Financial penalties alone in the commit-reveal scheme may not ensure the reliable generation of random numbers.
+Integrating VDFs with it, restructured as a commit-reveal-recover mechanism, can significantly enhance security.
 Figure 1: In our model, a user sends a VDF proof after the Pietrzak VDF evaluation.
@@
 Additionally, the actual implementation results were consistent with our theoretical expectations.
-  • Contrary to previous studies’ expectations, we demonstrated that the Pietrzak VDF could be utilized in the Ethereum environment with under 10 KB calldata size.
+  • Contrary to previous studies’ expectations, we demonstrated that the Pietrzak VDF could be utilized in the Ethereum environment with under 8 KB calldata size with the 2048-bit RSA key length.
 The rest of the paper is organized as follows: Section 2 introduces related studies about VDFs and their implementation.
@@
 Employing the Fiat-Shamir heuristic, both methods facilitate non-interactive verification, suitable for blockchain environments.
-However, the computational intensity and storage demands for managing large numeric values pose substantial hurdles.
+However, the computational intensity and memory demands for managing large numeric values pose substantial hurdles.
 Wesolowski’s VDF maintains a constant proof length ($O(1)$), independent of the time delay parameter $T$, making it particularly advantageous where proof size is a concern.
@@
 The experimental results from Attias et al.
-[3] show that a performance duration of approximately one minute at the 2048-bit level is feasible.
+[3] show that a performance duration of approximately one minute at the 2048-bit RSA key level is feasible.
 Consequently, we advocate for a similar time range in VDF applications.
@@
   • Zero byte: 4 gas
-Considering the above gas pricing, we can calculate the cost for sending one 2048-bit length data in the aforementioned data structure. 2048-bit data is transformed into 512-length hex letters.
+Considering the above gas pricing, we can calculate the cost for sending one 2048-bit length data in the aforementioned data structure.
+Taking into account the probability of each byte being non-zero, which is $\frac{1}{256}$, the expected gas cost for sending one 2048-bit length data is 1036 gas.
+The gas for the Big Number structure is added to this.
 In our implementation, the function verifyRecursiveHalvingProof processes the generalized verification (Algorithm 4). It is defined as the Listing 2 in Solidity codes.
@@
 For static types such as uint256, fields including delta and T are padded to 32 bytes.
-For dynamic types including bytes, struct, and dynamic arrays, they are encoded by specifying an offset to the start of its data area, followed by the length of the data and the actual data.
+On the other hand, In EVM, dynamic types, including bytes, struct, and dynamic arrays, are encoded by specifying an offset to the start of the data, its length, and then the data itself.
+This routine is applied recursively for nested dynamic types.
 In case of BigNumber[] memory v, it is the dynamic array of dynamic type struct BigNumber that contains dynamic bytes val.
 So the way dynamic types are encoded is recursively applied to each halving proof v.
-In EVM, dynamic types such as arrays and structs are encoded by specifying an offset to the start of the data, its length, and then the data itself.
-This method applies recursively for nested dynamic types.
 Parameters such as $T$, $\delta$, $\lambda$, and parameter count (e.g., 6 in our model) determine the offsets and lengths.
@@
 The intrinsic gas in EVM refers the minimum cost for a transaction.
-It contains the transaction cost and the cost for data allocation.
+It contains the constant transaction cost (21000 gas) and the cost for data supply.
 For verification, we need not only to allocate data in a transaction, but also to dispatch data into the verification function.
@@
 We name a naive exponentiation approach as exponentiation by squaring in the figures.
-Even if square and multiply (Algorithm 5 reduces the gas usage, it is obvious that the precompiled ModExp uses outstandingly less gas for the modular exponentiation than the others.
+The comparing data named ‘Precompile’ is the gas usage when only the precompiled ModExp is used for multiexponentiation.
+Even if square and multiply (Algorithm 5) reduces the gas usage, it is obvious that the precompiled ModExp uses outstandingly less gas for the modular exponentiation than the others.
 Figure 3(c) and Figure 3(d) show the gas usage experiment results for modular multiexponentiation.
 We implemented Dimitrov multiexponentitaion including the precompiled ModExp.
-The comparing data named ‘Precompile’ is the gas usage when only the precompiled ModExp is used for multiexponentiation.
 The latter takes significantly less gas than the Dimitrov algorithm.
@@
 Table 1 details the gas consumption and calldata sizes for the implemented parameters, uniformly using $\delta=9$.
-Contrary to the previous expectation of calldata sizes around 40KB [3], the sizes recorded were significantly lower, ranging from 5.52 to 7.35 KB for $\lambda=2048$ and from 7.22 to 9.72 KB for $\lambda=3072$.
-Gas usage varied from 1.96 to 2.80 million for $\lambda=2048$ and from 3.68 to 5.26 million for $\lambda=3072$.
+Contrary to the previous expectation of calldata sizes around 40KB [3], the sizes recorded were significantly lower, ranging from 5.38 to 7.25 KB for $\lambda=2048$ and from 7.13 to 9.63 KB for $\lambda=3072$.
+Gas usage varied from 1.97 to 2.80 million for $\lambda=2048$ and from 3.68 to 5.25 million for $\lambda=3072$.
 These findings not only demonstrate the efficacy of our approach but also highlight potential areas for further optimization in VDF implementations.
 Such improvements could lead to significant cost reductions in blockchain networks, enhancing the overall efficiency of cryptographic operations.
-Figure 5: Experiment results for halving proof skipping parameter $\delta$ with $\tau=22,23,24$
+Figure 5: Experiment results for halving proof skipping parameter $\delta$ with $\tau=20,21,22,23,24,25$
 (a) Skipping parts of halving proofs ($\lambda=2048$)
@@
 We anticipate that such an approach could be applicable not only within the EVM but also in more general computing environments.
-Although resource pricing in general systems would differ from the gas costs in the EVM, with elements like the repetition of halving protocols and resource consumption proportional to the proof length being predictable, the costs of exponentiation in VDF ($2^{2^{T}}$) are expected to be exponential.
+Although resource pricing in general systems would differ from the gas costs in the EVM, with elements like the repetition of halving protocols and resource consumption proportional to the proof length being predictable, the costs of exponentiation in VDF ($x^{2^{2^{\tau}}}$) are expected to be exponential.
 Therefore, based on the outcomes of this study, further research on optimizing Pietrzak VDFs in general computing environments appears reasonable.

### v4

@@
-Implementation Study of Cost-Effective Verification for Pietrzak’s Verifiable Delay Function in Ethereum Smart Contracts
+Implementation Study of Cost-Effective Verification for Pietrzak’s VDF in Ethereum Smart Contracts
 Abstract
 Verifiable Delay Function (VDF) is a cryptographic concept that ensures a minimum delay before output through sequential processing, which is resistant to parallel computing.
-Among the two well-known VDF protocols, Wesolowski and Pietrzak VDF, we focus on the Pietrzak VDF due to its computational efficiency and suitability for blockchain environments.
-Pietrzak’s approach uses a recursive proof verification with the halving protocol, offering a practical alternative despite the longer proof length than Wesolowski’s approach.
-Given the scarcity of research on practical VDF verification implementation, especially within smart contracts, this paper aims to implement cost-effective verification for the Pietrzak VDF in an Ethereum-based environment without compromising the VDF verification’s integrity and reliability.
-Firstly, we propose generalized proof generation and verification algorithms for potential efficiency improvement.
-Secondly, we categorize and measure the gas cost of each part in a transaction for VDF verification.
-Thirdly, based on the analysis, we theoretically predict the optimized proof construction.
-Finally, we demonstrate the theoretical prediction matches the implementation results.
-Furthermore, our research shows that the proof length of the Pietrzak VDF is generated under 8 KB with the 2048-bit RSA key length, much smaller than the previous expectation.
-This implies that the Pietrzak VDF can be practically used for cryptographic applications on blockchains.
+One of the significant VDF protocols academically reviewed is the VDF protocol proposed by Pietrzak.
+However, for the blockchain environment, the Pietrzak VDF has drawbacks including long poof size and recursive protocol computation.
+In this paper, we present an implementation study of Pietrzak VDF verification on Ethereum Virtual Machine (EVM).
+We found that the discussion in the Pietrzak’s original paper can help a clear optimization in EVM where the costs of computation are predefined as the specific amounts of gas.
+In our results, the cost of VDF verification can be reduced from 4M to 2M gas, and the proof length can be generated under 8 KB with the 2048-bit RSA key length, which is much smaller than the previous expectation.
 §§footnotetext: These authors contributed equally to this work.
@@
 1 Introduction
-footnotetext: We express our special thanks to Prof. Seungjoo Kim at Korea University for his invaluable advice on cryptographic algorithms.
-In blockchain technology, one critical element is the generation and use of cryptographically secure random numbers.
-These secure random numbers play a pivotal role in various applications of blockchain; for example, in consensus, random numbers can ensure fair selections of sequencers and block proposers based on a probabilistic process.
-Random numbers are essential for ensuring anonymity and serve as the basis for cryptographic techniques.
-Despite the significance of secure random numbers, incorporating them into smart contracts results in challenges.
-Due to the lack of random number methods in the smart contract interface, individuals previously used to apply future block data (i.e.
-the blockhash of a future block) as an ad-hoc random number.
-It can be manipulated by a block generator.
-Hence, leveraging ad-hoc random numbers is not a suitable approach for the present situation.
-In this context, two notable techniques have emerged: Verifiable Random Functions (VRFs) and Verifiable Delay Functions (VDFs).
-VRFs are cryptographic primitives that produce a random number along with proof of its correctness, allowing anyone to verify the number’s authenticity without compromising its unpredictability.
-Several commercial random number generating services using VRF, such as Chainlink VRF, are already available.
-As part of their operation, a node acting as an oracle transmits a randomly generated number from off-chain nodes to a blockchain.
-On the other hand, VDFs are designed to produce a unique output that requires a predetermined amount of time to compute, but whose result can be efficiently verified by others.
-Therefore, VDFs do not directly generate a random number.
-Rather, VDFs support random number generation of the commit-reveal scheme by utilizing their time-lock property.
-The most critical weakness of the commit-reveal scheme is that all the participants must reveal to generate a secure random number.
-VDFs’ time-lock property ensures that the random outcome cannot be predicted or accelerated by computational power in case of any absence.
-It makes VDFs practical for applications that benefit from a guaranteed delay in randomness generation.
-Financial penalties alone in the commit-reveal scheme may not ensure the reliable generation of random numbers.
-Integrating VDFs with it, restructured as a commit-reveal-recover mechanism, can significantly enhance security.
-Figure 1: In our model, a user sends a VDF proof after the Pietrzak VDF evaluation.
-The primary issue is the significant gas usage for the VDF verification.
-The secondary issue is the length of the VDF proof as the Pietrzak VDF’s length increases as the time delay increases.
-Pivotal advancements in VDFs have been made through the research of Pietrzak [1] and Wesolowski [2], who each proposed practical implementations of VDFs.
-Despite these theoretical developments, there remains a notable gap in research regarding the application of VDFs within blockchain environments.
-The challenge lies not only in the theoretical construction of VDFs, but also in adapting these constructions to the resource usage context of blockchain systems.
-For example, the Pietrzak VDF uses the halving protocol that makes the length of proof proportional to the delay time.
-It makes utilizing the Pietrzak VDF considered unrealistic in blockchain [3].
-The Wesolowski VDF needs a hash-to-prime oracle that is not supported in most blockchains.
-When implementing VDFs, the process can broadly be divided into three parts: evaluation, proof generation, and verification.
-In the context of blockchain, the verification component must be implemented on-chain.
-This segmentation is due to our belief that in typical blockchain applications, both evaluation and proof generation processes are performed off-chain, mainly due to their computational intensity and the associated costs when executed on a blockchain network.
-Our goal has been to efficiently implement VDF verification within smart contracts as Figure 1’s illustration.
-Specifically, we have targeted the Pietrzak VDF model.
-There are two reasons: Firstly, the Wesolowski VDF requires a hash-to-prime function that is challenging to implement on smart contracts.
-Secondly, we conjectured that optimization can mitigate the long-proof length issue partially.
-Based on this algorithm choice, consequently, this research demonstrated that the Pietrzak VDF verification can be utilized in the EVM-compatible smart contract environment through various cost-saving analysis and experiments.
+The Verifiable Delay Function (VDF)’s major role is to guarantee a specific delay for evaluation.
+It can play an important role in security applications.
+The most notable application of blockchain using VDF is a distributed randomness beacon.
+VDFs do not directly generate a random number.
+Rather, they help to guarantee the liveness of a random number generation scheme, Commit-Reveal, by recovering a random number when the reveal phase fails.
+This has led Ethereum, one of the largest cryptocurrencies, to include VDFs in their future roadmap for the RANDAO block proposer selection mechanism.
+In recent years, several VDF mechanisms have been proposed [1, 2, 3, 4]; however, the concept of VDFs remains relatively new.
+This indicates that we need comprehensive research on VDFs across various dimensions.
+For example, recently found parallel computation mechanisms for VDF evaluation [5, 6] show that we need more research on the squaring computation.
+On the other hand, there is a scarcity of implementation research [7] on VDFs.
+In this paper, we focus on VDF implementation on EVM, one of the leading blockchain environments.
+Our goal is to efficiently implement VDF verification within smart contracts, providing a foundation for EVM security applications utilizing VDFs.
+Specifically, we focus on the Pietrzak VDF [2] for two key reasons: First, the Wesolowski VDF [1], another major VDF, requires a hash-to-prime function, which is challenging to implement directly in smart contracts.
+Second, we hypothesize that the performance optimization technique discussed by Pietrzak [2] could significantly reduce the overall cost within the EVM environment.
+As a result of this focus, our research demonstrates through theoretical analysis and experiments that Pietrzak VDF verification can be effectively utilized within EVM-compatible smart contracts.
 The contributions of this paper are as follows:
   • To the best of our knowledge, this research presents the first academic study of Pietrzak VDF implementation on smart contracts.
-  • We proposed generalized algorithms for proof generation and verification of the Pietrzak VDF for potential efficiency enhancement.
   • Through theoretical analysis, we demonstrated the existence of a unique parameter that reduces the proof length.
@@
   • Contrary to previous studies’ expectations, we demonstrated that the Pietrzak VDF could be utilized in the Ethereum environment with under 8 KB calldata size with the 2048-bit RSA key length.
-The rest of the paper is organized as follows: Section 2 introduces related studies about VDFs and their implementation.
-Section 3 describes the proposed target algorithms and our implementation environment.
-Section 4 illustrates the data structures of proofs and analyzes gas costs for the data and its dispatching cost.
-Section 5 studies arithmetic operations in the verification process and their gas costs.
-In Section 6, based on the previous sections’ outcomes, we present theoretical analysis and the cost-effective implementation.
-Section 7 discusses this research’s contributions, limitations, and future directions.
-Finally, Section 8 concludes the paper.
-2 Background
-This section introduces VDFs, cryptographic algorithms crucial for blockchain applications and secure timestamping.
-It also reviews related works that inform the VDF implementations discussed in this paper.
-2.1 Verifiable Delay Functions
-VDFs are cryptographic protocols crafted to ensure that their execution takes a predefined minimum amount of time, regardless of the computational power applied.
-Here we introduce the core mechanisms that define VDFs and the core properties that VDFs get from the elements.
-2.1.1 Definition and Fundamental Properties
-A VDF consists of three key algorithms that facilitate one round: setup, evaluation, and verification.
-  • Setup $(\text{Setup}(1^{\lambda})\rightarrow(\mathbb{G},T))$: Initializes the VDF with public parameters.
-Here, $1^{\lambda}$ represents the security parameter indicating the required security level, $\mathbb{G}$ is a cryptographic group used in computations, and $T$ is the time delay parameter that sets the duration of the computation.
-  • Evaluation $(\text{Evaluate}(\mathbb{G},T,X)\rightarrow(Y,\pi))$: For a given input $X$, the VDF computes an output $Y$ over time $T$, which is deliberately designed to be sequential, preventing acceleration through parallel processing.
-The output includes a proof $\pi$ that validates the computation.
-  • Verification $(\text{Verify}(\mathbb{G},T,X,Y,\pi)\rightarrow\{\textsf{true},\textsf{false}\})$: Enables any observer to quickly and efficiently verify the correctness of the output $Y$ and its proof $\pi$ without undergoing the time-intensive computation themselves.
-A return of true confirms the validity of the output, while false indicates an error.
-From the above algorithms, VDFs must get the three primary attributes [4]:
-  • Sequentiality in evaluation: The computation must take a fixed minimum time, irrespective of the available computational resources.
-  • Efficiency in verification: Although the computation itself is intentionally delayed, the verification process is designed to be quick and require minimal resources.
-  • Uniqueness in output: Each input to a VDF yields a unique, unpredictable output, ensuring reliability and consistency in applications.
-2.2 Feasibility of On-Chain VDF Verification
-Both Wesolowski [2] and Pietrzak [1] VDFs operate by transforming a value $x$ into $y=x^{2^{T}}\mod N$, where $N$ is an RSA modulus, although their verification methods differ.
-Scalability and implementation costs are significant challenges in blockchain systems.
-Employing the Fiat-Shamir heuristic, both methods facilitate non-interactive verification, suitable for blockchain environments.
-However, the computational intensity and memory demands for managing large numeric values pose substantial hurdles.
-Wesolowski’s VDF maintains a constant proof length ($O(1)$), independent of the time delay parameter $T$, making it particularly advantageous where proof size is a concern.
-The implementation on blockchain platforms like Ethereum is particularly challenging due to the high cost associated with verifying or finding primes for large numbers, such as 2048 bits.
-This necessity significantly increases the implementation cost, making full on-chain execution unfeasible.
-Pietrzak’s VDF does not require operations like hash-to-prime, simplifying its on-chain implementation.
-The protocol involves a halving process, where $T$ is divided by two until it becomes one, reducing the complexity of computations with each iteration.
-However, the proof length and cost are influenced by the time delay variable $T$, and the proof length grows with $O(\log_{2}(T))$.
-This characteristic can result in verifications that exceed the size limit for feasible blockchain implementation.
-The research anticipated that the length of the proof will approximate 40 KB [3].
-2.3 Related Works
-In this subsection, we review the critical advancements in VDFs, from their theoretical foundations to practical implementations related to this study.
-Theoretical research on VDFs.
-The concept of VDFs has garnered significant attention for its potential to enable new applications in cryptography and blockchain technology.
-The foundational work by Rivest, Shamir, and Wagner on "Time-Lock Puzzles and Timed-Release Crypto" [5] laid the groundwork for the development of protocols that secure data until a certain amount of time has passed.
-This seminal paper introduced the concept of time-lock puzzles, a precursor to VDFs, demonstrating how cryptographic operations could be used to create a delay mechanism that is practically infeasible to bypass without solving the puzzle.
-Building upon these principles, Pietrzak [1] and Wesolowski [2] independently proposed two distinct approaches to VDFs, each with unique advantages in terms of efficiency and security.
-Pietrzak’s method focuses on using iterated sequential squaring in a group of unknown order, providing a balance between computation time and verification efficiency.
-Wesolowski’s approach, on the other hand, introduces an elegant solution that reduces the verification complexity, making it more practical for applications that require fast and efficient verification processes.
-Further advancing the field, Dan Boneh et al.
-[6] have conducted extensive research on optimizing VDFs, addressing critical aspects such as the construction of more efficient protocols and the reduction of computational requirements for both generation and verification phases.
-Their work not only highlights the theoretical advancements in VDF technology but also emphasizes the practical implications for secure, decentralized systems.
-Implementation of VDFs.
-The practical realization of VDFs has become an area of intense research and development, seeking to bridge the gap between theoretical robustness and real-world applicability.
-A significant contribution to this endeavor is the study by Attias et al.
+The rest of the paper is organized as follows: Section 2 reviews related works on VDFs.
+Section 3 describes the target algorithm and experiment environment.
+In Section 4, we measure the gas costs for VDF verification on the EVM and formulate them into a function using regression.
+Section 5 analyzes the optimal parameters for VDF verification costs and presents experimental results that demonstrate the theoretical analysis.
+Finally, Section 6 concludes the paper with a discussion of further research directions.
+The source code for our implementation is available in the Appendix A.
+2 Related Works
+The concept of recursive computation for delayed output was firstly proposed as the "Time-Lock Puzzle" by Rivest et al.
+[8].
+The Time-Lock Puzzle does not include efficient public verifiability of the solution.
+This was later addressed by Mahmoody et al.
+in their Proof of Sequential Work (PoSW) [9].
+Lenstra and Wesolowski proposed Sloth, which is considered the first Verifiable Delay Function (VDF) style mechanism [3].
+Boneh et al.
+[10] formalized the concept of VDF.
+Following the introduction of VDFs, various mechanisms were proposed, among which the VDFs by Wesolowski [1] and Pietrzak [2] had unique advantages in terms of efficiency and security.
+For the use in blockchain, Veedo and Minroot [4] were proposed for Starknet and RANDAO respectively.
+However, recent findings by Biryukov et al.
+[5] and Leurent [6] demonstrate that distributed algorithms for modular squaring exist.
+This implies that the fundamental assumption of using sequential squaring puzzles is not valid, highlighting the need for further research on this issue.
+Not only theoretical research, but also implementation studies are crucial for approaching practical applications.
+However, such studies appear to have been conducted in a rather limited scope.
+A significant contribution is the study by Attias et al.
 [7], which provides a comprehensive implementation analysis of two leading VDFs, Pietrzak and Wesolowski.
@@
 Choi et al.
-[7] proposed three practical commit-reveal-recover techniques.
-Their protocols are crucial for establishing the commit-reveal-recover scheme, which mostly existed at the conceptual level.
+[11] proposed three practical commit-reveal-recover protocols.
 They incorporated the Wesolowski VDF into Ethereum smart contracts for the purpose of demonstration.
 Our research covers VDF verification, which corresponds to the Proof-of-Exponentiation (PoE) stated in their work.
-Their PoE gas usage is approximately 2.3 million, which is comparable to the gas consumption of our work with the time delay variable of $2^{22}$.
-Due to the Wesolowski VDF not having increased calculation requirements with greater time delays, our approach becomes less efficient when the time delay variable exceeds $2^{23}$.
-Nevertheless, the Wesolowski VDF encounters an issue when it comes to fully implementing the hash-to-prime function as previously indicated.
 3 Implementation Target and Environment
 In this section, we provide the target algorithm and environment.
-Firstly, we define the Pietrzak VDF algorithms we aim to implement in this research.
-Secondly, we describe the experiment environment in the following sections.
-3.1 Pietrzak VDF Algorithms
-Our research has a model with a smart contract that supports a function that performs Pietrzak VDF verification.
-Referring to the Pietrzak’s original paper [1], we define the evaluation as described in Algorithm 1 and the halving protocol for proof generation in Algorithm 2.
-3.1.1 VDF Evaluation
-$y\equiv x^{2^{T}}\equiv x^{2^{2^{\tau}}}\mod{N}$ (1)
-The VDF evaluation process as defined in Algorithm 1 is pivotal in generating a VDF output from given inputs.
-It basically computes Equation 1.
-The algorithm takes an initial value $x$, a time parameter $T$ (also notated as $2^{\tau}$), and a modulus $N$ to produce the output $y$.
-$N$ that is a $\lambda$-bit number is produced by an RSA setup ($N=p\cdot q$ where $p,q$ are two big prime numbers).
-The process involves exponentiating $x$ to $2^{T}\mod N$ to ensure the difficulty and time consumption of the evaluation are as intended.
-This evaluation serves as the foundation for verifying that the output can only be computed with a predefined amount of real-time computational work, which is crucial in applications requiring guarantees against parallelization.
-Algorithm 1 Evaluation of the Pietrzak VDF
-1:input: $x,T,N$
-2:output: $y$
-3:$y\leftarrow x$
-4:for $k\leftarrow 1$ to $T$ do
-5:     $y\leftarrow y^{2}\mod N$
-6:end for
-7:return $y$
-3.1.2 Halving Protocol
-The halving protocol detailed in Algorithm 2 optimizes the proof generation process for VDFs by reducing the verification time incrementally.
-At each step, the protocol halves the time parameter $T$, recalculates intermediate values $x_{i}$ and $y_{i}$, and updates these values based on a cryptographic hash function to ensure integrity.
-As we implement it on Ethereum, we use Keccak256 as the cryptographic hash function.
-This iterative reduction continues until the desired granularity is reached, ensuring that the verifier’s workload is minimized while maintaining cryptographic security.
-Algorithm 2 Halving Protocol of the Pietrzak VDF
-1:input: $N,x_{i},T,y_{i}$
-2:output: $N,x_{i+1},T/{2^{i}},y_{i+1},\mu_{i+1}$
-3:$\mu_{i}\leftarrow x_{i}^{2^{T/{2^{i}}}}$
-4:$r_{i}\leftarrow\text{Hash}(x_{i}+y_{i}+v_{i})$
-5:$x_{i+1}\leftarrow x_{i}^{r_{i}}\cdot v_{i}\mod N$
-6:if $T/2^{i-1}$ is odd then
-7:     $y_{i}\leftarrow y_{i}^{2}\mod N$
-8:end if
-9:$y_{i+1}\leftarrow v_{i}^{r_{i}}\cdot y_{i}\mod N$
-10:return $N,x_{i+1},T/{2^{i}},y_{i+1},\mu_{i+1}$
-3.1.3 Introduction of Generalized Proof Generation/Verification and Its Rationale
-In the original paper [1] and the related works [6, 3], the conventional Pietrzak halving protocol is repeated until the time parameter $T$ gets 1 for proof generation.
-It helps the verifier consume the nearly least resources to verify VDF proofs in general systems.
-However, on Ethereum, the cost of resource usage, called gas, significantly differs from general environments.
-Ethereum Yellowpaper [8] provides predefined gas prices for transactions and opcodes executions.
-Especially in this context, repeating the halving protocol till $T=1$ does not guarantee the most efficient verification.
-Therefore, we generalized the Pietrzak verification algorithm with flexible repetition of the halving protocol.
-In Algorithm 4, we introduced a new parameter $\delta$.
-The halving protocol is repeated till the time variable becomes equal to $\delta$.
-Then, the verifier checks $y\stackrel{{\scriptstyle?}}{{=}}x^{2^{2^{\delta}}}$.
-3.1.4 Generalized Proof Generation
-In Algorithm 3, the generalized proof generation is used to construct the cryptographic proofs required for VDF verification under flexible conditions.
-The proof generation starts with the evaluation result and halves the exponentiation of a proof step-by-step until reaching the specified parameter $\delta$.
-Each halving can be tracked as intermediate values $v_{i}$ are stored as part of the proof.
-This mechanism allows for flexible adjustment of the verification process to accommodate different system capacities and security requirements.
-Algorithm 3 Generalized Proof Generation of the Pietrzak VDF with Evaluation
-1:input: $x,T,\delta,N$
-2:output: $\{\pi_{i}\}_{i=1}^{\tau-\delta}$
+Firstly, we define the Pietrzak VDF verification algorithm we aim to implement.
+Secondly, we describe the experiment environment for the cost measurement.
+3.1 Pietrzak VDF verification
+Algorithm 1 outlines the Pietrzak VDF verification process.
+Given that a prover insists the VDF evaluation $x^{2^{T}}=y$ with the time delay parameter $T=2^{\tau}$, the set of halving proof $\{\pi_{i}\}_{i=1}^{\tau}$, and the modulus $N$.
+It repeats the halving protocol (lines 5 – 13) until the verifier can simply check the result by squaring the base.
+The halving protocol reduces the number of exponentiations required for verification by half.
+The halving protocol outputs $x_{1},y_{1}$ and the proof $\pi$ reasoning $x_{1}^{2^{2^{\tau-1}}}=y_{1}$.
+Therefore, the verifier essentially needs to repeat the halving protocol $\tau$ times for the minimal exponentiation.
+One drawback of the Pietrzak VDF is that as the time delay parameter increases, the proof size and the number of halving protocol repetitions for the verifier also increase.
+In the original paper proposing the Pietrzak VDF, a method to enhance efficiency was suggested.
+Since the halving protocol also consumes resources, the verifier can repeat the halving protocol $\tau-\delta$ times instead of $\tau$ times.
+Then, the verifier needs to check $x^{2^{2^{\delta}}}\stackrel{{\scriptstyle?}}{{=}}y$.
+Furthermore, this approach helps to reduce the size of the proof.
+Algorithm 2 describes this refined verification procedure.
+In most systems, the costs of each computation and the overhead due to proof size are dependent on the specific system environment, making it difficult to determine the optimal value for the proof shortening parameter, $\delta$.
+However, in the EVM, the costs for data (calldata) and arithmetic computations are predefined.
+This led us to hypothesize that it is possible to identify the optimal $\delta$ value for implementation, thereby minimizing both computation and calldata costs in blockchain environments.
+Algorithm 1 Pietrzak VDF Verification
+1:input: $x,y,T,\{\pi_{i}\}_{i=1}^{\tau},N$
+2:output: True or False
 3:$\tau\leftarrow\lfloor\log_{2}(T)\rfloor$
-4:$y\leftarrow x$
-5:for $k\leftarrow 1$ to $T$ do
-6:     $y\leftarrow y^{2}\mod N$
-7:end for
-8:$(x_{1},y_{1})\leftarrow(x,y)$
-9:for $i\leftarrow 1$ to $\tau-\delta$ do
-10:     $v_{i}\leftarrow x^{2^{T/2^{i}}}$
-11:     $r_{i}\leftarrow keccak256(x_{i}+y_{i}+v_{i})$
-12:     $x_{i+1}\leftarrow x_{i}^{r_{i}}\cdot v_{i}\mod N$
-13:     if $T/2^{i-1}$ is odd then
-14:         $y_{i}\leftarrow y_{i}^{2}\mod N$
-15:     end if
-16:     $y_{i+1}\leftarrow v_{i}^{r_{i}}\cdot y_{i}\mod N$
-17:     $\pi_{i}\leftarrow v_{i}$
-18:end for
-19:return $\{\pi_{i}\}_{i=1}^{\tau-\delta}$
-3.1.5 Generalized Verification
-Algorithm 4 describes the generalized verification process for VDFs on Ethereum.
-It utilizes the proofs generated by the generalized proof generation method to validate the sequence of operations leading to the final result.
-The verifier iteratively computes and checks each step of the proof against the corresponding intermediate values and the final condition $y_{\tau-\delta+1}=x_{\tau-\delta+1}^{2^{2^{\delta}}}$.
-This process ensures that the computation was performed correctly and meets the specified delay constraints, which is crucial for applications relying on the integrity of timed cryptographic operations.
-Algorithm 4 Generalized Verification of the Pietrzak VDF
+4:$(x_{1},y_{1})\leftarrow(x,y)$
+5:for $i\leftarrow 1$ to $\tau$ do
+6:     $v_{i}\leftarrow\pi_{i}$
+7:     $r_{i}\leftarrow keccak256(x_{i}||y_{i}||v_{i})$
+8:     $x_{i+1}\leftarrow x_{i}^{r_{i}}\cdot v_{i}\mod N$
+9:     if $T/2^{i-1}$ is odd then
+10:         $y_{i}\leftarrow y_{i}^{2}\mod N$
+11:     end if
+12:     $y_{i+1}\leftarrow v_{i}^{r_{i}}\cdot y_{i}\mod N$
+13:end for
+14:if $y_{\tau+1}=x_{\tau+1}^{2}$ then
+15:     return True
+16:else
+17:     return False
+18:end if
+Algorithm 2 Refined Pietrzak VDF Verification
 1:input: $x,y,T,\delta,\{\pi_{i}\}_{i=1}^{\tau-\delta},N$
@@
 6:     $v_{i}\leftarrow\pi_{i}$
-7:     $r_{i}\leftarrow keccak256(x_{i}+y_{i}+v_{i})$
+7:     $r_{i}\leftarrow keccak256(x_{i}||y_{i}||v_{i})$
 8:     $x_{i+1}\leftarrow x_{i}^{r_{i}}\cdot v_{i}\mod N$
@@
 18:end if
-According to Algorithm 3, and 4, the data should be submitted are the root ($x$), the VDF evaluation result ($y=x^{2^{T}}$), the time delay variable ($T$), the VDF proof ($\{\pi_{i}\}_{i=1}^{\tau-\delta}$), and the group description ($N$) as described in Equation 2.
-$\Big{(}N,x,y,\tau,\overbrace{(\pi_{1},\pi_{2},\pi_{3},...,\pi_{k})}^{\overbrace{\lfloor\log_{2}(T)\rfloor}^{\tau}-\delta}\Big{)}$ (2)
-3.2 Target Time Delay and the Bit Length
-Recent advances in VDF implementation using C++, as studied by Attias et al.
-[3], demonstrate that VDF evaluation times can range from $2^{20}$ to $2^{25}$ cycles.
-These findings suggest that the implementations are not only effective but also efficient in experimental setups.
-In the context of Ethereum, where block generation typically requires around 12 seconds and even less time in other blockchain systems, we recommend that the time delay for generating VDF-based random numbers should not exceed one minute.
-This recommendation aims to prevent excessive resource consumption without compromising practical utility.
-The experimental results from Attias et al.
-[3] show that a performance duration of approximately one minute at the 2048-bit RSA key level is feasible.
-Consequently, we advocate for a similar time range in VDF applications.
-The security of our target Pietrzak VDF configuration depends on the difficulty of the prime factoring problem, akin to that in the RSA encryption system.
+3.2 Target Bit Length and Target Delay
+The security of Pietrzak VDF configuration depends on the difficulty of the prime factoring problem, akin to that in the RSA encryption system.
 Currently, a 2048-bit key length is deemed sufficient for maintaining security standards.
@@
 Therefore, our experiments included tests with both 2048-bit and 3072-bit settings for VDFs.
+In a recent study on VDF implementation using C++ [7], VDF evaluation times with the time delay parameter $T$ ranging from $2^{20}$ to $2^{25}$ can be completed within one minute with a personal CPU.
+The evaluation involves performing $y=x^{2^{T}}$.
+In other words, the target range for $\tau$ is 20 – 25 where $T=2^{\tau}$.
+We chose to conduct our VDF experiments within the same parameter range, as it strikes a balance between computational feasibility and practical utility while avoiding excessive resource consumption.
 3.3 Environment for Gas Usage Measurement
-Here we describe the setup of this research used to measure gas usage in Ethereum smart contracts, detailing the use of Solidity and the related framework.
-3.3.1 Development Environment and Compiler Version
-The development of the contract was conducted using Solidity version 0.8.23.
-This choice was driven by the compatibility and full support provided by the Hardhat framework, a comprehensive Ethereum development environment with integrated debugging capabilities.
-At the start of the development process, Solidity 0.8.23 was the latest compiler version fully supported by Hardhat for debugging and other essential development functionalities.
-Generally, gas optimization is best achieved with the latest version of the compiler.
+3.3.1 Development Environment and Tools
+The development of the contract was conducted using Solidity version 0.8.26 and EVM version Cancun.
+At the start of the development process, it was the latest compiler version fully supported by Foundry, which is a comprehensive Ethereum development toolchain that provides a robust environment for building, testing, and debugging smart contracts.
+Furthermore, gas optimization is best achieved with the latest version of the compiler.
 3.3.2 Optimizer Configuration
-For configuration, the core parameter of the Solidity optimizer is Runs which indicates how often each opcode will be used in the deployed contract.
+The core parameter for the Solidity optimizer configuration is Runs which indicates how often each opcode will be used in the deployed contract.
 That is, higher Runs implies less gas cost for execution and higher gas cost for deployment.
@@
 Instead, via-IR processes an intermediate step where Yul code, an intermediate representation, is generated.
-In this step, we can enhance the Yul optimizer’s efficiency.
-We replaced the msize() operation with the free memory pointer.
+In this step, we could enhance the Yul optimizer’s efficiency by replacing the msize() operation with the free memory pointer.
 This adjustment necessitated manual updates to the memory indexes, a process rigorously tested to ensure the integrity of our implementation.
@@
 This change not only aligns with Solidity’s memory model but also activates Via-IR in the compiler settings, resulting in significant gas savings.
-3.3.3 Gas Measurement
-We implemented external functions in MinimalApplication.sol, the major source file for implementation, to measure the gas used with different strategies.
-The gas cost depends on the order of the function selectors in these external functions, but the differences are considered to be negligible, with a maximum difference of about 60 gas used across the seven external functions.
-Therefore, an error margin of about 60 gas exists for the results presented in the figures.
-3.3.4 Local Testing Environment
-We utilized the Hardhat node, provided by the Hardhat framework, to create a local testing environment.
-The amount of gas usage was obtained using the eth_estimateGas method, which generates and returns an estimate of the gas required to complete the transaction.
-This method accurately measures the gas used in the execution of different contract functions and provides valuable insights into the efficiency of the implementations.
+3.3.3 Gas Measurement in Local Testing Enviornment
+As the test target code, we implemented contracts for each test in one main source file ensuring that each contract contains only one external function.
+This approach minimizes the error margin that can arise when multiple external functions are present, as the order of the function selectors impacts the gasUsed.
+For the cost (gas usage) measurement, we utilized the Anvil node, provided by the Foundry toolchain, to create a local testing environment.
+The gas usage was obtained using the Forge testing framework provided by Foundry.
+The Forge standard library’s Test contract includes the lastCallGas() function, which retrieves the gasUsed in the last call.
+Additionally, isolate flag must be set when running the Forge’s test command to ensure all top-level calls are executed as separate transactions in distinct EVM contexts.
+This setup allows for more precise gas accounting.
 It is important to note that the actual gas cost incurred on the real network is determined by multiplying the estimated gasUsed by the gasPrice at the Ethereum network circumstance.
-4 Data and Accompanying Gas Cost
+4 Cost measurement and approximation
 In this section, we introduce the data structure we used to send VDF proofs in the Ethereum transaction calldata.
@@
 However, the EVM natively supports arithmetic operators only up to 256-bit numbers.
-Therefore, to facilitate efficient computation of numbers exceeding 256 bits, it is necessary to employ customized data formats.
+To facilitate efficient computation of numbers exceeding 256 bits, it is necessary to employ customized data formats.
 We revised the big number library implementation for Solidity developed by Firo [12].
@@
 Listing 1 Calldata for a Big Number Example
-4.2 Gas Cost of Calldata
-The gas cost for Ethereum transaction calldata was fixed in EIP-2028 "Transaction data gas cost reduction" [10]. According to the document, the cost of calldata is divided into two cases:
-  • Non-zero byte: 16 gas
-  • Zero byte: 4 gas
+4.2 Gas Cost for Calldata and Dispatch
+The gas cost for Ethereum transaction calldata, as described in EIP-2028 "Transaction data gas cost reduction" [13], is divided into two cases: 16 gas per non-zero byte and 4 gas per zero byte.
 Considering the above gas pricing, we can calculate the cost for sending one 2048-bit length data in the aforementioned data structure.
@@
 The gas for the Big Number structure is added to this.
-In our implementation, the function verifyRecursiveHalvingProof processes the generalized verification (Algorithm 4). It is defined as the Listing 2 in Solidity codes.
-1function verifyRecursiveHalvingProof(
-2 BigNumber[] memory v,
-3 BigNumber memory x,
-4 BigNumber memory y,
-5 BigNumber memory n,
-6 uint256 delta,
-7 uint256 T
-8)
-Listing 2 Interface of the Halving Verification Function
+In our implementation, the function verifyRecursiveHalvingProof processes the refined verification, Algorithm 2.
 A call to this function involves encoding parameter values.
-Start by computing the keccak256 of the UTF-8 byte representation of the string111verifyRecursiveHalvingProof((bytes,uint256)[],(bytes,uint256),(bytes,uint256),(bytes,uint256),uint256,uint256) based on Listing 2 and use the first four bytes, which results in 0xd8e6ac60.
+The process starts by computing the keccak256 of the UTF-8 byte representation of the string111verifyRecursiveHalvingProof((bytes,uint256)[],(bytes,uint256),(bytes,uint256),(bytes,uint256),uint256,uint256), then using the first four bytes, which results in 0xd8e6ac60.
 For static types such as uint256, fields including delta and T are padded to 32 bytes.
@@
 Finally, we can predict the approximate gas usage of calldata using this EVM data construction structure.
-Considering the precision of data and its cost, we performed experiments to test the size of calldata and the related gas usage.
+Considering the prediction of data and its cost, we performed experiments to test the size of calldata and the related gas usage.
 The related gas usage comprises of the intrinsic gas and data dispatching gas.
@@
 Therefore, we decided to include the dispatching cost into the data cost in this study.
-Figure 2(a) for $\lambda=2048$ and 2(b) for $\lambda=3072$ show the calldata size with the number of proofs and the related gas usage.
+Figure 1(a) for $\lambda=2048$ and 1(b) for $\lambda=3072$ show the calldata size with the number of proofs and the related gas usage.
 We can see the gas usage and the calldata size simply and linearly increase with the number of proofs.
-Figure 2: Experiment results of the calldata length and cost for the proofs
-(a) Calldata size and gas cost including intrinsic gas and dispatching gas ($\lambda=2048$)
-(b) Calldata size and gas cost including intrinsic gas and dispatching gas ($\lambda=3072$)
-Finally, to elucidate the relationship between transaction size and gas costs, we perform a regression analysis.
-By applying the least squares method to approximate the best fit, we derive the following linear regression equations 3 and 4 for data sizes of 2048 and 3072 bytes, respectively:
+Figure 1: Experiment results of the calldata and its dispatch cost
+(a) $\lambda=2048$
+(b) $\lambda=3072$
+Figure 2: Experiment results of the halving algorithm cost
+(a) $\lambda=2048$
+(b) $\lambda=3072$
+Figure 3: Experiment results of the ModExp cost
+(a) $\lambda=2048$
+(b) $\lambda=3072$
+Finally, to describe the relationship between transaction size and gas costs, we performed a basic regression analysis.
+By applying the least squares method to approximate the best fit, we derive the following linear regression equations 1 and 2 for data sizes of 2048 and 3072 bytes, respectively:
 $\mathcal{G}_{data,2048}(\tau)=5295.73\cdot\tau+37610.11$ (1)
@@
 These equations estimate the gas cost $\mathcal{G}$ as a function of transaction size $\tau$, highlighting the increase in required gas with larger transaction calldata.
-5 Arithmetic Operations and Accompanying Gas Cost
-This section focuses on analyzing the expenses associated with performing arithmetic operations in the context of verifying the Pietrzak VDF.
-The cost consists of two components: the cost of repeated halving and the cost of modular exponentiation for the final check.
-The algorithm executes a halving operation and then proceeds with a modular exponentiation verification.
-The modular exponentiation cost is studied before the halving cost, as the halving process includes the modular exponentiation.
-Using the results, we perform a regression analysis to formulate mathematical equations for the cost in the section that follows.
 4.3 Modular Exponentiation Cost
-This part explores strategies for optimizing the computational cost of modular exponentiation, a frequent operation in various cryptographic algorithms.
-It discusses algorithms such as Square and Multiply [11], Dimitrov multiexponentiation [12], and the ModExp contract introduced by EIP-198.
-5.1.1 Square and Multiply
-The square and multiply algorithm, also called binary exponentiation, is an efficient approach for computing exponentiation operations $x^{n}$, particularly beneficial for large number modular exponentiation.
-As described in Algorithm 5, by converting the exponent $b$ into its binary representation, the algorithm optimizes computation through a series of square (squaring) and multiply (multiplying by the base $x$) operations based on the binary digits of $n$.
-For an exponent represented by $k$ bits, the algorithm requires at most $k-1$ squaring and $k-1$ multiplication operations, making the computational complexity $O(\log_{2}n)$.
-This optimization is simple but significantly enhances efficiency for large-scale modular exponentiation tasks.
-Algorithm 5 Square and Multiply for Modular Exponentiation
-1:input: $a,b,N$
-2:output: $a^{b}$ mod $N$
-3:$r\leftarrow 1$
-4:$a\leftarrow a\mod N$
-5:while $b>0$ do
-6:     if $b\mod 2=1$ then
-7:         $r\leftarrow(r\cdot a)\mod N$
-8:     end if
-9:     $b\leftarrow b>>1$
-10:     $a\leftarrow(a\cdot a)\mod N$
-11:end while
-12:return $r$
-5.1.2 Dimitrov Multiexponentiation
-The Dimitrov algorithm [12, 3] presented in Algorithm 6 enhances the efficiency of large integer multiplication, crucial in fields such as cryptography and computational mathematics, by significantly reducing the computational complexity.
-Traditionally, the multiplication of large integers ($x^{a}y^{b}$) exhibits a complexity of $O(2\log_{2}max(a,b))$.
-Utilizing the binary exponentiation method [13], the Dimitrov algorithm achieves a computational complexity closer to $O(\frac{7}{4}\log_{2}max(a,b))$.
-This characteristic allows for the rapid and resource-efficient multiplication of very large integers, making the algorithm particularly valuable for applications requiring efficient large-scale numerical operations.
-In the generalized verification (Algorithm 4), lines 9 – 12 perform multiexponentiation only if $T/2$ is odd.
-Therefore, we considered the Dimitrov algorithm as a candidate algorithm to apply for optimization.
-Algorithm 6 Dimitrov Multiexponentiation Algorithm
-1:input: $x,y,a,b,N$
-2:output: $x^{a}\cdot y^{b}\mod N$
-3:$h\leftarrow\max(\lfloor\log_{2}a\rfloor+1,\lfloor\log_{2}b\rfloor+1)$
-4:$z\leftarrow 1$
-5:$q\leftarrow x\cdot y\mod N$
-6:for $i=h-1$ down to 00 do
-7:     $z\leftarrow z\cdot z\mod N$
-8:     if $a_{i}=1$ and $b_{i}=0$ then
-9:         $z\leftarrow z\cdot x\mod N$
-10:     else if $a_{i}=0$ and $b_{i}=1$ then
-11:         $z\leftarrow z\cdot y\mod N$
-12:     else if $a_{i}=1$ and $b_{i}=1$ then
-13:         $z\leftarrow z\cdot q\mod N$
-14:     end if
-15:end for
-16:return $z$
-Figure 3: Experiment results of the modular exponentiation cost
-(a) Gas usage of exponentiation by squaring, square and multiply, and precompiled ModExp, x is 2048-bit level
-(b) Gas usage of exponentiation by squaring, square and multiply, and precompiled ModExp, x is 3072-bit level
-(c) Gas usage of Dimitrov multiexponentiation and only precompiled ModExp, x is 2048-bit level
-(d) Gas usage of Dimitrov multiexponentiation and only precompiled ModExp, x is 3072-bit level
-5.1.3 ModExp by EIP-198
-EIP-198 [14], known for introducing "Big Integer Modular Exponentiation", introduces the ModExp (Modular Exponential) precompiled contract to the EVM, offering a streamlined and cost-efficient approach to conducting large integer modular exponentiation, a foundational operation in numerous cryptographic algorithms.
-This function calculates $a^{b}\mod n$, where $a$, $b$, and $n$ are large numbers, and prior to EIP-198, such execution on-chain was not officially supported by any operations.
-The ModExp contract mitigates these expenses by providing a standardized computation method, thus facilitating the incorporation of cryptographic computations within smart contracts.
-Building on this, EIP-2565 [15] contributed to further refining the gas cost model for these modular exponentiation operations by proposing adjustments to the pricing formula.
-5.1.4 Comparison of Exponentiation Techniques and Regression
-Figure 3(a) and Figure 3(b) present the gas usage experiment results of three modular exponentiation techniques.
-We name a naive exponentiation approach as exponentiation by squaring in the figures.
-The comparing data named ‘Precompile’ is the gas usage when only the precompiled ModExp is used for multiexponentiation.
-Even if square and multiply (Algorithm 5) reduces the gas usage, it is obvious that the precompiled ModExp uses outstandingly less gas for the modular exponentiation than the others.
-Figure 3(c) and Figure 3(d) show the gas usage experiment results for modular multiexponentiation.
-We implemented Dimitrov multiexponentitaion including the precompiled ModExp.
-The latter takes significantly less gas than the Dimitrov algorithm.
-Therefore, we applied ModExp for both the modular exponentiation and multiexponentiation operations in the implementation in Algorithm 4.
-In the experiment, the pricing of ModExp is linearly proportional to the number of exponentiation as $\mathcal{G}(T)=aT+b$.
-As Equation 1, we denote $\tau$ as the logarithmic value of T.
+EIP-198 [14], known for "Big Integer Modular Exponentiation", introduced the ModExp (Modular Exponential) precompiled contract to the EVM, offering a streamlined and cost-efficient approach for large integer modular exponentiation.
+[^2] EIP-2565 [15] contributed to further refining the gas cost model for these modular exponentiation operations by proposing adjustments to the pricing formula.
+As its obvious that a precompile provides the best efficinecy, we applied ModExp for both the modular exponentiation and multi-exponentiation operations in the implementation in Algorithm 2.
+In Figure 3(a) and Figure 3(b) illustrating the experiment results, the pricing of ModExp is linearly proportional to the number of exponentiation as $\mathcal{G}(T)=aT+b$, where we denote $\tau$ as the logarithmic value of $T$.
 We could get the following results from a simple regression process.
@@
 The halving cost of the arithmetic operations in the verification indicates the gas usage for lines 5 – 13 in Algorithm 2.
-The modular exponentiation operations in lines 8, 10, and 12 use the precompiled ModExp since its best efficiency in the previous subsection.
+The modular exponentiation operations in lines 8, 10, and 12 use the precompiled ModExp.
 For simplicity, as we assumed $T=2^{\tau}$, in this case, the if statement in line 9 is always false, therefore line 10 is never executed.
-Hence, the gas usage will be clearly proportional to the number of the repetition time, that is, $\tau-\delta$.
-Figure 4: Experiment results of the halving repetition cost
-(a) Gas usage of the halving algorithm ($\lambda=2048$)
-(b) Gas usage of the halving algorithm ($\lambda=3072$)
+Hence, the gas usage is clearly proportional to the number of the repetition time, that is, $\tau-\delta$.
 The results of our experiments demonstrate the anticipated correlation.
@@
 The data exhibits a clear linear relationship, allowing us to perform a regression analysis and model the results using a linear function.
-By employing the method of least squares approximation, we obtain the subsequent functions as Equation 7 and 8:
-$\mathcal{G}_{halving,2048}(\tau)=165482.26\cdot\tau-155642.3$ (7)
-$\mathcal{G}_{halving,3072}(x)=306089.32\cdot\tau-238572.65$ (8)
+By employing the method of least squares approximation, we obtain the subsequent functions as Equation 5 and 6:
+$\mathcal{G}_{halving,2048}(\tau)=160604.41\cdot\tau-157178.66$ (5)
+$\mathcal{G}_{halving,3072}(x)=299896.98\cdot\tau-248119.67$ (6)
 5 Cost-Effective Implementation
@@
 5.1 Theoretical Analysis on Proof Generation for Gas Usage Optimization
-The optimized proof generation in this research means minimizing the verification gas cost. We can define the total gas cost function $g_{total}$ using the previous regression results.
+The optimized proof generation in this research means minimizing the verification gas cost in EVM. We can define the total gas cost function $g_{total}$ using the previous regression results.
 $\mathcal{G}_{total}(\tau,\delta)=\mathcal{G}_{calldata}+\mathcal{G}_{halving}+\mathcal{G}_{exp}+C$ (7)
@@
 Proof.
-Firstly, we calculate the partial derivative of $\mathcal{G}_{total}$ with respect to $\delta$, which yields
-$\frac{\partial\mathcal{G}_{total}}{\partial\delta}=\ln{2}\cdot\gamma\cdot 2^{\delta}-(\alpha+\beta).$
-To show that there is a unique $\delta$ that minimizes $\mathcal{G}_{total}$, we need to find where this derivative equals zero. This leads us to the equation
-$2^{\delta}=\frac{\alpha+\beta}{\ln{2}\cdot\gamma}.$
-The function $2^{\delta}$ is strictly increasing. Hence, the equation $2^{\delta}=\frac{\alpha+\beta}{\ln{2}\cdot\gamma}$ has a unique solution $\delta$, which can be expressed as
-$\delta=\log_{2}(\alpha+\beta)-\log_{2}(\ln{2}\cdot\gamma).$
-This $\delta$ is where the first derivative equals zero.
-Further, to confirm that this $\delta$ is indeed a minimum, we need to check the second derivative.
-The monotonic nature of the exponential function in this context implies that as $\delta$ increases, $\gamma\cdot 2^{\delta}$ grows without bounds, while decreasing $\delta$ below this point leads to $\gamma\cdot 2^{\delta}$ becoming smaller than $\alpha+\beta$, thereby confirming that the function is minimized at this $\delta$.
-Thus, $\mathcal{G}_{total}(\tau,\delta)$ indeed has one unique minimum at $\delta$ within the specified range of $\tau>\delta\geq 0$, completing the proof. ∎
+See Appendix B. ∎
 Also, we denote the minimizing $\delta$ as
@@
 From Theorem 1, we can get the unique $\delta_{m}$ optimizing the gas cost.
-Applying all the regression results in Equation 3, 4, 5, 6, 7, and 8, we get the optimizing $\delta_{m}\approx 9.47$ for $\lambda=2048$ and $\delta_{m}\approx 9.18$ for $\lambda=3072$ respectively.
+Applying all the regression results in Equation 1, 2, 3, 4, 5, and 6, we get the optimizing $\delta_{m}\approx 9.43$ for $\lambda=2048$ and $\delta_{m}\approx 9.15$ for $\lambda=3072$ respectively.
 In practice, $\delta$ for implementation must be an integer.
@@
     If $\Delta>-\log_{2}(\ln 2)$, then $\mathcal{G}(\tau,\lfloor\delta_{m}\rfloor)>\mathcal{G}(\tau,\lceil\delta_{m}\rceil)$.
+Proof.
+See Appendix C. ∎
+Since $\log_{2}(\ln 2)\approx-0.52$, the optimal integer value of $\delta$ is 9 for both $\lambda=2048$ and $3072$, as the 1st case of Corollary 2.
+5.2 Optimized Verification Gas Cost
+Figures 4(a) and 4(b) depict the total gas cost for Pietrzak VDF verification at $\lambda=2048$ and $\lambda=3072$ respectively.
+Consistent with the theoretical predictions in Theorem 1, the gas usage graph shows a decline to a minimal point before increasing sharply.
+Optimal gas costs were observed at $\delta=9$ for both configurations, aligning perfectly with the findings of Corollary 2.
+This alignment between theoretical expectations and empirical results confirms the accuracy of our model and highlights its potential for optimizing Pietrzak VDF verification in blockchain applications.
+[…]
+Figure 4: Results for halving proof skipping parameter $\delta$ with $\tau=20,21,22,23,24,25$ for $y=x^{2^{2^{\tau}}}$
+(a) Skipping parts of halving proofs ($\lambda=2048$)
+(b) Skipping parts of halving proofs ($\lambda=3072$)
+Table 1: Implementation results based on the analysis
+$\lambda$ $\delta$ $\tau$ Gas used Calldata size (KB)
+2048 9 20 1817855 5.38
+21 1973222 5.75
+22 2129116 6.13
+23 2286058 6.50
+24 2445803 6.88
+25 2603753 7.25
+3072 9 20 3495396 7.13
+21 3776893 7.63
+22 4078468 8.13
+23 4362630 8.63
+24 4658937 9.13
+25 4958138 9.63
+6 Conlusion and Future Works
+Our study demonstrates that Pietrzak’s VDF can be effectively optimized for the EVM, reducing gas costs from 4M to 2M and minimizing proof size to under 8 KB with a 2048-bit RSA key.
+These optimizations validate the theoretical models and show practical viability.
+While our paper does not address the Wesolowski VDF implementation, which poses challenges due to its hash-to-prime function, exploring indirect or approximate implementation methods could make it feasible for the EVM.
+Future work should focus on utilizing Wesolowski VDF verification on EVM as it has a big benefit on the constant and short proof length and verification cost.
+We plan to implement on-chain Wesolowski VDF verification both directly and indirectly.
+For indirect implementation, a hash-to-prime function could be executed in a Truebit-style proof-of-fraud system.
+For direct implementation, refining the hash-to-prime function to be viable and secure on the EVM is essential.
+Appendix A Code availability
+To facilitate reproducibility and peer review, we have made the source code for our implementation available anonymously.
+The repository contains all the scripts and supplementary materials necessary to reproduce the results discussed in this paper.
+You can access the repository at the following link: https://github.com/usgeeus/Pietrzak-VDF-solidity-verifier
+Appendix B Proof of Theorem 1
+Proof.
+Firstly, we calculate the partial derivative of $\mathcal{G}_{total}$ with respect to $\delta$, which yields
+$\frac{\partial\mathcal{G}_{total}}{\partial\delta}=\ln{2}\cdot\gamma\cdot 2^{\delta}-(\alpha+\beta).$
+To show that there is a unique $\delta$ that minimizes $\mathcal{G}_{total}$, we need to find where this derivative equals zero. This leads us to the equation
+$2^{\delta}=\frac{\alpha+\beta}{\ln{2}\cdot\gamma}.$
+The function $2^{\delta}$ is strictly increasing. Hence, the equation $2^{\delta}=\frac{\alpha+\beta}{\ln{2}\cdot\gamma}$ has a unique solution $\delta$, which can be expressed as
+$\delta=\log_{2}(\alpha+\beta)-\log_{2}(\ln{2}\cdot\gamma).$
+This $\delta$ is where the first derivative equals zero.
+Further, to confirm that this $\delta$ is indeed a minimum, we need to check the second derivative.
+The monotonic nature of the exponential function in this context implies that as $\delta$ increases, $\gamma\cdot 2^{\delta}$ grows without bounds, while decreasing $\delta$ below this point leads to $\gamma\cdot 2^{\delta}$ becoming smaller than $\alpha+\beta$, thereby confirming that the function is minimized at this $\delta$.
+Thus, $\mathcal{G}_{total}(\tau,\delta)$ indeed has one unique minimum at $\delta$ within the specified range of $\tau>\delta\geq 0$, completing the proof. ∎
+Appendix C Proof of Corollary 2
 Proof.
@@
 In the same way, we can get the other two cases. ∎
-Since $\log_{2}(\ln 2)\approx-0.52$, the optimal integer value of $\delta$ is 9 for both $\lambda=2048$ and $3072$, as the 1st case of Corollary 2.
-6.2 Optimized Verification Gas Cost
-Figures 5(a) and 5(b) depict the total gas cost for Pietrzak VDF verification at $\lambda=2048$ and $\lambda=3072$ respectively.
-Consistent with the theoretical predictions in Theorem 1, the gas usage graph shows a decline to a minimal point before increasing sharply.
-Optimal gas costs were observed at $\delta=9$ for both configurations, aligning perfectly with the findings of Corollary 2.
-This alignment between theoretical expectations and empirical results confirms the accuracy of our model and highlights its potential for optimizing Pietrzak VDF verification in blockchain applications.
-Table 1 details the gas consumption and calldata sizes for the implemented parameters, uniformly using $\delta=9$.
-Contrary to the previous expectation of calldata sizes around 40KB [3], the sizes recorded were significantly lower, ranging from 5.38 to 7.25 KB for $\lambda=2048$ and from 7.13 to 9.63 KB for $\lambda=3072$.
-Gas usage varied from 1.97 to 2.80 million for $\lambda=2048$ and from 3.68 to 5.25 million for $\lambda=3072$.
-These findings not only demonstrate the efficacy of our approach but also highlight potential areas for further optimization in VDF implementations.
-Such improvements could lead to significant cost reductions in blockchain networks, enhancing the overall efficiency of cryptographic operations.
-Figure 5: Experiment results for halving proof skipping parameter $\delta$ with $\tau=20,21,22,23,24,25$
-(a) Skipping parts of halving proofs ($\lambda=2048$)
-(b) Skipping parts of halving proofs ($\lambda=3072$)
-Table 1: Implementation results based on the analysis
-$\lambda$ $\delta$ $\tau$ Gas used Calldata size (KB)
-2048 9 20 1970844 5.38
-21 2133372 5.75
-22 2297629 6.13
-23 2470594 6.50
-24 2620907 6.89
-25 2799805 7.25
-3072 9 20 3680363 7.13
-21 3967114 7.63
-22 4318050 8.13
-23 4614955 8.63
-24 4947040 9.13
-25 5254521 9.63
-7 Discussion
-This section discusses the implementation of VDFs on Ethereum, focusing on programming choices for Ethereum’s Virtual Machine and their impact on gas optimization.
-We introduce challenges associated with VDF deployment on Ethereum’s Layer 1 due to high costs and explore Layer 2 solutions.
-The discussion concludes with an analysis of economic factors critical for VDF integration and sustainability on the blockchain.
-7.1 Implementation Languages
-There are main development languages in EVM-compatible network environments, for example, Viper, Huff, Solidity, and Yul.
-For our implementation, which requires dynamically adjusting proof lengths, we excluded Vyper due to its inability to handle dynamic data structures efficiently.
-Instead, we chose Solidity combined with Yul, leveraging Solidity for its comprehensive features and developer familiarity, and Yul for its superior gas optimization capabilities.
-This combination provides a balance between high-level language safety and low-level control over the Ethereum Virtual Machine.
-Considering future enhancements, we are exploring Huff, noted for its exceptional gas efficiency and direct control over the EVM.
-Huff originated from the Aztec team where contracts needed extreme efficiency, even more than what Yul could provide.
-Although Huff offers significant optimization opportunities, it lacks the built-in safety checks of higher-level languages, raising the implementation complexity.
-Nevertheless, given our implementation’s emphasis on gas optimization, the benefits of adopting Huff justify its consideration for future development.
-7.2 Practicality of VDF on EVM
-Implementing VDFs on Ethereum’s Layer 1 presents significant challenges primarily due to high gas costs and the computational intensity required.
-Our detailed optimization trials reveal that despite employing cutting-edge optimization techniques, verifying a Pietrzak VDF on Layer 1 still demands at least one million gas units.
-This high resource consumption makes VDF applications impractical for widespread use on Ethereum’s primary network, especially for operations requiring frequent and repetitive VDF verifications.
-The economic implications of these requirements can stifle innovation and discourage developers from incorporating VDFs into their decentralized applications, limiting the exploration of potentially valuable use cases such as randomized governance protocols or secure public randomness in blockchain networks.
-However, the emergence of Layer 2 solutions, such as optimistic rollups and zkRollups, offers a promising alternative for leveraging VDFs within the Ethereum ecosystem.
-These technologies provide a secondary processing layer that can handle complex computations off the main Ethereum chain, thereby significantly reducing gas costs and improving transaction throughput.
-Platforms like Optimism capitalize on these advantages, offering gas prices that are often thousands of times cheaper than those on Layer 1.
-This dramatic cost reduction opens up new possibilities for using VDFs on Ethereum, allowing for more frequent and economically viable implementations.
-Layer 2 solutions not only mitigate the high cost barriers associated with Layer 1 but also broaden the scope for adopting VDFs in a variety of applications, thus encouraging further development and research into this promising area of blockchain technology.
-7.3 Lack of Economic VDF Protocols in Ethereum
-The integration of VDFs into smart contracts holds great promise for enhancing blockchain security and reliability.
-However, beyond the technical challenges, there exists a substantial hurdle in the lack of a robust economic model specifically designed to support the unique operational needs of VDFs, such as incentivizing both provers and verifiers and managing commit-reveal-recover protocols.
-The absence of such an economic framework complicates the practical deployment of VDFs as it does not adequately address compensation for participants who dedicate substantial computational resources to generate and verify proofs.
-This gap can deter the participation of potential contributors, ultimately hindering the scalability and effectiveness of VDFs in real-world blockchain applications.
-Moreover, considering the decentralized essence of blockchain technology, it is crucial that any protocol, especially those as intricate as VDFs, is underpinned by an economic strategy that not only supports technical execution but also ensures economic sustainability without the need for central governance.
-This calls for innovative research into economic models that are specifically tailored to the decentralized context of blockchain networks.
-These models need to effectively balance incentives and deterrence, fostering a system where honest participation is rewarded while dishonest behavior is penalized.
-Such economic strategies are vital for maintaining the integrity and fairness of the blockchain ecosystem, ensuring that VDFs are not only technically feasible but also economically viable and beneficial to all participants, thereby aligning with the overarching principles of decentralization and trust minimization in blockchain systems.
-7.4 Future Works on a General Environment
-Through this research, we investigated the efficient implementation of Pietrzak VDF verification within the Ethereum Virtual Machine (EVM) environment and observed results that align both theoretically and experimentally.
-During our study, we identified significant effects of generalized proof generation and verification, which were also included within the predictable realms of theoretical analysis.
-We anticipate that such an approach could be applicable not only within the EVM but also in more general computing environments.
-Although resource pricing in general systems would differ from the gas costs in the EVM, with elements like the repetition of halving protocols and resource consumption proportional to the proof length being predictable, the costs of exponentiation in VDF ($x^{2^{2^{\tau}}}$) are expected to be exponential.
-Therefore, based on the outcomes of this study, further research on optimizing Pietrzak VDFs in general computing environments appears reasonable.
-8 Conclusion
-In conclusion, our research has successfully demonstrated the feasibility and significance of implementing VDFs within blockchain environments, particularly through smart contracts in Ethereum.
-By focusing on the Pietrzak VDF construction, we have not only bridged a crucial gap in the academic studies of VDFs but also highlighted their practical utility in enhancing the security and efficiency of blockchain applications.
-Contrary to previous expectations, our work represents the first academic result in the cost-effective on-chain verification of the Pietrzak VDF, showing significant potential in ensuring cryptographically secure random number generation across various blockchain applications.
-Through strategic implementation, we have achieved substantial cost reductions, making the Pietrzak VDF more accessible and applicable.
-This research opens new avenues for the integration of VDFs in blockchain systems, promising to enhance security, fairness, and functionality in the evolving distributed ledger technologies.

### v5

@@
 One of the significant VDF protocols academically reviewed is the VDF protocol proposed by Pietrzak.
-However, for the blockchain environment, the Pietrzak VDF has drawbacks including long poof size and recursive protocol computation.
+However, for the blockchain environment, the Pietrzak VDF has drawbacks including long proof size and recursive protocol computation.
 In this paper, we present an implementation study of Pietrzak VDF verification on Ethereum Virtual Machine (EVM).
@@
 Our goal is to efficiently implement VDF verification within smart contracts, providing a foundation for EVM security applications utilizing VDFs.
-Specifically, we focus on the Pietrzak VDF [2] for two key reasons: First, the Wesolowski VDF [1], another major VDF, requires a hash-to-prime function, which is challenging to implement directly in smart contracts.
+Specifically, we focus on the Pietrzak VDF [2] for two key reasons: First, the Wesolowski VDF [1], another major VDF, requires a hash-to-prime function, which is challenging to implement directly in smart contracts (see B).
 Second, we hypothesize that the performance optimization technique discussed by Pietrzak [2] could significantly reduce the overall cost within the EVM environment.
@@
   • To the best of our knowledge, this research presents the first academic study of Pietrzak VDF implementation on smart contracts.
-  • Through theoretical analysis, we demonstrated the existence of a unique parameter that reduces the proof length.
-Additionally, the actual implementation results were consistent with our theoretical expectations.
-  • Contrary to previous studies’ expectations, we demonstrated that the Pietrzak VDF could be utilized in the Ethereum environment with under 8 KB calldata size with the 2048-bit RSA key length.
+  • Through theoretical analysis, we demonstrated the existence of a unique parameter that reduces the proof length and the execution gas fee, making the implementation more cost-efficient.
+  • We demonstrated that the Pietrzak VDF could be utilized in the Ethereum environment with under 8 KB calldata size for a 2048-bit RSA key length.
+This calldata size represents the serialized data required for VDF verification and highlights the feasibility of implementing the Pietrzak VDF on Ethereum smart contracts, contrary to previous studies’ expectations.
 The rest of the paper is organized as follows: Section 2 reviews related works on VDFs.
@@
 Section 5 analyzes the optimal parameters for VDF verification costs and presents experimental results that demonstrate the theoretical analysis.
-Finally, Section 6 concludes the paper with a discussion of further research directions.
-The source code for our implementation is available in the Appendix A.
+Section 6 addresses challenges in implementing VDFs on Ethereum and explores potential directions for research and optimization.
+Finally, Section 7 concludes the paper by summarizing our findings and outlining future research plans.
+The source code for our implementation is available in A.
 2 Related Works
-The concept of recursive computation for delayed output was firstly proposed as the "Time-Lock Puzzle" by Rivest et al.
-[8].
-The Time-Lock Puzzle does not include efficient public verifiability of the solution.
-This was later addressed by Mahmoody et al.
-in their Proof of Sequential Work (PoSW) [9].
-Lenstra and Wesolowski proposed Sloth, which is considered the first Verifiable Delay Function (VDF) style mechanism [3].
-Boneh et al.
-[10] formalized the concept of VDF.
-Following the introduction of VDFs, various mechanisms were proposed, among which the VDFs by Wesolowski [1] and Pietrzak [2] had unique advantages in terms of efficiency and security.
-For the use in blockchain, Veedo and Minroot [4] were proposed for Starknet and RANDAO respectively.
-However, recent findings by Biryukov et al.
-[5] and Leurent [6] demonstrate that distributed algorithms for modular squaring exist.
-This implies that the fundamental assumption of using sequential squaring puzzles is not valid, highlighting the need for further research on this issue.
-Not only theoretical research, but also implementation studies are crucial for approaching practical applications.
-However, such studies appear to have been conducted in a rather limited scope.
-A significant contribution is the study by Attias et al.
-[7], which provides a comprehensive implementation analysis of two leading VDFs, Pietrzak and Wesolowski.
-Their work meticulously evaluates the performance of VDFs’ evaluation, proof generation, and verification.
-Their work provided us with insight into the research and optimization techniques for VDFs.
-However, they expected the Pietrzak VDF would not be viable for blockchains.
-This is contrary to our results that presented the feasibility of the Pietrzak VDF on Ethereum smart contracts.
+Decentralized, bias-resistant randomness is a cornerstone of blockchain security, driving extensive research into Distributed Randomness Beacon (DRB) protocols.
+Recent Systematization of Knowledge (SoK) studies by Raikwar and Gligoroski [8], Choi et al.
+[9], and Kavousi et al.
+[10] provide comprehensive taxonomies of DRBs, analyzing them in terms of cryptographic primitives, interactivity, unbiasability, unpredictability, and scalability.
+These surveys underscore the critical role of publicly verifiable randomness in preventing adversarial bias or manipulation in decentralized applications.
+One cryptographic mechanism that has gained significant traction in DRB protocols is the Verifiable Delay Function (VDF).
+The concept of VDFs builds on earlier advances in time-delayed cryptography.
+Rivest et al.
+introduced the “Time-Lock Puzzle” [11], which envisioned recursive computation for delayed output but lacked efficient public verifiability.
+Mahmoody et al.
+later proposed Proof of Sequential Work (PoSW) [12], thereby formalizing sequential tasks in cryptographic protocols.
+Lenstra and Wesolowski’s Sloth [3] is often regarded as a precursor to modern VDFs, which were later formalized by Boneh et al.
+[13] into efficient and publicly verifiable delay primitives.
+Following the initial formalization of VDFs, various constructions have emerged.
+Wesolowski’s [1] and Pietrzak’s [2] designs are particularly notable for their strong security proofs and computational efficiency.
+Beyond DRBs, VDFs have also found applications in non-interactive cryptographic timestamping [14] and DoS mitigation [15], underscoring their versatility.
+Pietrzak’s VDF, in particular, has been adapted in several ways: Hoffmann and Pietrzak [16] extend it to watermarkable and zero-knowledge settings, Motepalli and Jacobsen [17] apply it to blockchain bootstrapping, and it used the Pietrzak implementation by Chia Network [18].
+However, these works focus primarily on conceptual or protocol-level enhancements, rather than on optimizing on-chain performance in environments like the Ethereum Virtual Machine (EVM).
+In addition to these foundational VDF-based designs, other DRB solutions have been proposed that rely on various cryptographic primitives to achieve robust randomness on-chain.
+For example, RandRunner [19] employs trapdoor VDFs to ensure strong uniqueness, Bicorn [20] incorporates VDFs in an optimistically efficient DRB, and RandChain [21] leverages sequential Proof-of-Work to decentralize randomness generation.
+Raikwar [22] further examines competitive DRB protocols, highlighting different strategies for achieving resilience against adversarial manipulation.
+Collectively, these systems reinforce the importance of secure, verifiable primitives that balance theoretical soundness with practical feasibility in blockchain contexts.
+Later proposals, such as Veedo and Minroot [4], target specific blockchain ecosystems (e.g., Starknet or RANDAO).
+At the same time, recent research by Biryukov et al.
+[5] and Leurent [6] indicates that distributed modular squaring might be possible, challenging the assumption of a strictly sequential computation.
+This highlights the need to reconcile theoretical security assumptions with real-world performance constraints in VDF-based DRBs.
+Despite increasing interest in VDFs, implementation-focused investigations remain sparse.
+One of the few in-depth evaluations is by Attias et al.
+[7], who examined both Pietrzak’s and Wesolowski’s VDFs—covering evaluation, proof generation, and verification—and noted that large proof sizes could be problematic for blockchain environments.
 Choi et al.
-[11] proposed three practical commit-reveal-recover protocols.
-They incorporated the Wesolowski VDF into Ethereum smart contracts for the purpose of demonstration.
-Our research covers VDF verification, which corresponds to the Proof-of-Exponentiation (PoE) stated in their work.
+[20] later integrated Wesolowski’s VDF into Ethereum commit-reveal schemes but did not provide a cost-optimization analysis or a large-scale feasibility study.
+To address this gap, our work examines the on-chain performance of Pietrzak’s VDF on Ethereum smart contracts.
+We concentrate on gas costs, proof sizes, and optimization strategies that render VDF-based solutions viable.
+By offering a comprehensive feasibility analysis for Pietrzak’s VDF in real-world blockchain environments, we extend existing DRB research and demonstrate that VDFs can be efficiently deployed within the EVM.
 3 Implementation Target and Environment
 In this section, we provide the target algorithm and environment.
-Firstly, we define the Pietrzak VDF verification algorithm we aim to implement.
+Firstly, we define the Pietrzak VDF algorithm we aim to implement.
 Secondly, we describe the experiment environment for the cost measurement.
-3.1 Pietrzak VDF verification
+3.1 Pietrzak VDF algorithms
+Our research has a model with a smart contract that supports a function that performs Pietrzak VDF verification.
+Referring to the Pietrzak’s original paper [2], we present the evaluation procedure and the halving protocol.
+In the following subsections, we show how these procedures lead to proof generation and verification in a blockchain environment, culminating in a refined approach that reduces on-chain overhead in the Ethereum Virtual Machine (EVM).
+3.1.1 VDF Evaluation
+The Pietrzak VDF algorithm operates with the following parameters:
+  • $x$: The input value for the VDF evaluation.
+  • $N$: The modulus (an RSA modulus $N=p\cdot q$), where $p$ and $q$ are large primes.
+  • $T$: The time delay parameter, where $T=2^{\tau}$ for some $\tau$.
+  • $y$: The output value, computed as $y=x^{2^{T}}\bmod N$.
+To evaluate the VDF, one exponentiates $x$ by $2^{T}\bmod N$, as shown in Equation 1:
+$y\equiv x^{2^{T}}\equiv x^{2^{2^{\tau}}}\mod{N}.$ (1)
+The sequential nature of repeated squaring ensures that this operation cannot be significantly parallelized, making it an effective time-delay mechanism.
+Algorithm 1 details this process, serving as the foundation for proving that the output $y$ can only be computed by incurring a predefined amount of real-time work.
+This property is crucial for applications requiring verifiable latency and resistance to parallel computation.
+Algorithm 1 Evaluation of the Pietrzak VDF
+1:input: $x,T,N$
+2:output: $y$
+3:$y\leftarrow x$
+4:for $k\leftarrow 1$ to $T$ do
+5:     $y\leftarrow y^{2}\mod N$
+6:end for
+7:return $y$
+3.1.2 Pietrzak VDF proof generation
+The proof generation defined in Algorithm 2 starts with the evaluation result and is used to construct the cryptographic proofs required for VDF verification.
+The halving protocol detailed in lines 5 – 13 optimizes the proof generation process for VDFs by reducing the verification time incrementally.
+At each step, the protocol halves the time parameter $T$, recalculates intermediate values $x_{i}$ and $y_{i}$, and updates these values based on a cryptographic hash function to ensure integrity.
+As we implement it on Ethereum, we use Keccak256 as the cryptographic hash function.
+This iterative reduction continues until the desired granularity is reached, ensuring that the verifier’s workload is minimized while maintaining cryptographic security.
+Each halving can be tracked as intermediate values $v_{i}$ are stored as part of the proof.
+Algorithm 2 Pietrzak VDF proof generation
+1:input: $x,T,N$
+2:output: $\{\pi_{i}\}_{i=1}^{\tau}$
+3:$\tau\leftarrow\lfloor\log_{2}(T)\rfloor$
+4:$(x_{1},y_{1})\leftarrow(x,y)$
+5:for $i\leftarrow 1$ to $\tau$ do
+6:     $v_{i}\leftarrow x^{2^{T/2^{i-1}}}$
+7:     $r_{i}\leftarrow keccak256(x_{i}||y_{i}||v_{i})$
+8:     $x_{i+1}\leftarrow x_{i}^{r_{i}}\cdot v_{i}\mod N$
+9:     if $T/2^{i-1}$ is odd then
+10:         $y_{i}\leftarrow y_{i}^{2}\mod N$
+11:     end if
+12:     $y_{i+1}\leftarrow v_{i}^{r_{i}}\cdot y_{i}\mod N$
+13:     $\pi_{i}\leftarrow v_{i}$
+14:end for
+15:return $\{\pi_{i}\}_{i=1}^{\tau}$
+3.1.3 Pietrzak VDF verification
 Algorithm 3 outlines the Pietrzak VDF verification process.
@@
 Therefore, the verifier essentially needs to repeat the halving protocol $\tau$ times for the minimal exponentiation.
-One drawback of the Pietrzak VDF is that as the time delay parameter increases, the proof size and the number of halving protocol repetitions for the verifier also increase.
-In the original paper proposing the Pietrzak VDF, a method to enhance efficiency was suggested.
-Since the halving protocol also consumes resources, the verifier can repeat the halving protocol $\tau-\delta$ times instead of $\tau$ times.
-Then, the verifier needs to check $x^{2^{2^{\delta}}}\stackrel{{\scriptstyle?}}{{=}}y$.
-Furthermore, this approach helps to reduce the size of the proof.
+One drawback of the Pietrzak VDF is that as the time delay parameter $T$ increases, the proof size and the number of halving protocol repetitions for the verifier also grow.
+To address this, Pietrzak introduced a proof-shortening parameter, $\delta$, which allows the verifier to skip $\delta$ halving rounds.
+Specifically, the verifier can repeat the halving protocol $\tau-\delta$ times instead of $\tau$ times, then check $x^{2^{2^{\delta}}}\stackrel{{\scriptstyle?}}{{=}}y$.
+The choice of $\delta$ may be determined by the system designer, the verifier, or specified by protocol parameters, depending on the implementation.
+This method both reduces the proof size and alleviates verification overhead.
 Algorithm 4 describes this refined verification procedure.
@@
 18:end if
+Figure 1: Algorithms for Pietrzak VDF Verification and Refined Verification.
 3.2 Target Bit Length and Target Delay
@@
 Therefore, our experiments included tests with both 2048-bit and 3072-bit settings for VDFs.
-In a recent study on VDF implementation using C++ [7], VDF evaluation times with the time delay parameter $T$ ranging from $2^{20}$ to $2^{25}$ can be completed within one minute with a personal CPU.
+In a recent study on VDF implementation using C++ [7], VDF evaluation times with the time delay parameter $T=2^{\tau}$ were examined.
+The study found that for $\tau$ values ranging from 20 to 25, the evaluation can be completed within one minute using a personal CPU.
 The evaluation involves performing $y=x^{2^{T}}$.
-In other words, the target range for $\tau$ is 20 – 25 where $T=2^{\tau}$.
-We chose to conduct our VDF experiments within the same parameter range, as it strikes a balance between computational feasibility and practical utility while avoiding excessive resource consumption.
+We chose to conduct our VDF experiments within the same $\tau$ range, as it strikes a balance between computational feasibility and practical utility while avoiding excessive resource consumption.
 3.3 Environment for Gas Usage Measurement
 3.3.1 Development Environment and Tools
-The development of the contract was conducted using Solidity version 0.8.26 and EVM version Cancun.
+The development of the smart contract was conducted using Solidity version 0.8.26 and EVM version Cancun.
 At the start of the development process, it was the latest compiler version fully supported by Foundry, which is a comprehensive Ethereum development toolchain that provides a robust environment for building, testing, and debugging smart contracts.
@@
 We set the optimizer parameter Runs to 4,294,967,295($2^{32}-1$), which is the maximum value of the parameter, indicating the highest level of optimization for execution.
-We set up Via-IR (Intermediate Representation) and the Yul optimizer in the Solidity compiler to generate efficient bytecode.
-Via-IR helps efficiency, security, and simplicity in code generation by deviating from the direct compilation of Solidity code to EVM bytecode.
-Instead, via-IR processes an intermediate step where Yul code, an intermediate representation, is generated.
-In this step, we could enhance the Yul optimizer’s efficiency by replacing the msize() operation with the free memory pointer.
-This adjustment necessitated manual updates to the memory indexes, a process rigorously tested to ensure the integrity of our implementation.
-We also applied memory-safe assembly blocks in our code.
-This change not only aligns with Solidity’s memory model but also activates Via-IR in the compiler settings, resulting in significant gas savings.
-3.3.3 Gas Measurement in Local Testing Enviornment
+We set up viaIR (Intermediate Representation) and the Yul optimizer in the Solidity compiler to generate efficient bytecode.
+viaIR helps efficiency, security, and simplicity in code generation by deviating from the direct compilation of Solidity code to EVM bytecode.
+Instead, viaIR processes an intermediate step where Yul code, an intermediate representation, is generated.
+The Yul optimizer refines this intermediate representation, enhancing the IR produced by the pipeline, as well as inline assembly and utility Yul code generated by the compiler [23].
+The msize() operation is incompatible with the Yul optimizer, necessitating its replacement with the free memory pointer.
+This adjustment required manual updates to memory indexes, which were rigorously tested to ensure the integrity of our implementation.
+To comply with Solidity’s memory model and enable viaIR, we incorporated memory-safe assembly blocks.
+These blocks ensure robust memory management and adhere to Solidity’s safety constraints, allowing the viaIR pipeline to be activated, resulting in significant gas savings [24].
+3.3.3 Gas Measurement in Local Testing Environment
 As the test target code, we implemented contracts for each test in one main source file ensuring that each contract contains only one external function.
@@
 Also, it utilizes the bit length data field to optimize multiplication and comparison operations.
-Therefore, we format the bytes data and pre-calculate bit length off-chain before submission, significantly reducing unnecessary on-chain computations.
+Therefore, we format the bytes data and include the bit length alongside it, significantly reducing unnecessary on-chain computations.
 For example, the data format for a 2048-bit representation in JSON is as shown in Listing 1.
@@
 Finally, we can predict the approximate gas usage of calldata using this EVM data construction structure.
-Considering the prediction of data and its cost, we performed experiments to test the size of calldata and the related gas usage.
+To estimate the gas usage for VDF proof calldata, we predicted the data composition and its associated costs by analyzing the probability distribution of zero and non-zero bytes in the serialized data.
+This prediction was derived using the EVM’s gas pricing model for calldata, where zero bytes incur 4 gas units and non-zero bytes incur 16 gas units as defined in EIP-2028 [26].
+Based on these predictions, we conducted experiments to measure the actual calldata size and the corresponding gas usage under various proof configurations.
 The related gas usage comprises of the intrinsic gas and data dispatching gas.
@@
 Therefore, we decided to include the dispatching cost into the data cost in this study.
-Figure 1(a) for $\lambda=2048$ and 1(b) for $\lambda=3072$ show the calldata size with the number of proofs and the related gas usage.
+Figure 2(a) for $\lambda=2048$ and 2(b) for $\lambda=3072$ show the calldata size with the number of proofs and the related gas usage.
 We can see the gas usage and the calldata size simply and linearly increase with the number of proofs.
-Figure 1: Experiment results of the calldata and its dispatch cost
-(a) $\lambda=2048$
-(b) $\lambda=3072$
-Figure 2: Experiment results of the halving algorithm cost
-(a) $\lambda=2048$
-(b) $\lambda=3072$
-Figure 3: Experiment results of the ModExp cost
-(a) $\lambda=2048$
-(b) $\lambda=3072$
+(a) Calldata Cost, $\lambda=2048$
+(b) Calldata Cost, $\lambda=3072$
+(c) Halving Cost, $\lambda=2048$
+(d) Halving Cost, $\lambda=3072$
+(e) ModExp Cost, $\lambda=2048$
+(f) ModExp Cost, $\lambda=3072$
+Figure 2: Experiment results: calldata cost, halving cost, and ModExp cost for different configurations ($\lambda=2048$ and $\lambda=3072$).
 Finally, to describe the relationship between transaction size and gas costs, we performed a basic regression analysis.
-By applying the least squares method to approximate the best fit, we derive the following linear regression equations 1 and 2 for data sizes of 2048 and 3072 bytes, respectively:
+By applying the least squares method [27], which finds the best-fit line by minimizing the sum of the squared differences between observed and predicted values, we derive the following linear regression equations 2 and 3 for data sizes of 2048 and 3072 bytes, respectively:
 $\mathcal{G}_{data,2048}(\tau)=5295.73\cdot\tau+37610.11$ (2)
@@
 As its obvious that a precompile provides the best efficinecy, we applied ModExp for both the modular exponentiation and multi-exponentiation operations in the implementation in Algorithm 4.
-In Figure 3(a) and Figure 3(b) illustrating the experiment results, the pricing of ModExp is linearly proportional to the number of exponentiation as $\mathcal{G}(T)=aT+b$, where we denote $\tau$ as the logarithmic value of $T$.
+In Figure 2(e) and Figure 2(f) illustrating the experiment results, the pricing of ModExp is linearly proportional to the number of exponentiation as $\mathcal{G}(T)=aT+b$, where we denote $\tau$ as the logarithmic value of $T$.
 We could get the following results from a simple regression process.
@@
 The results of our experiments demonstrate the anticipated correlation.
-Figure 2(a) and Figure 2(b) display the gas cost of the repeated halving operations with $\lambda=2048$ and $\lambda=3072$ correspondingly.
+Figure 2(c) and Figure 2(d) display the gas cost of the repeated halving operations with $\lambda=2048$ and $\lambda=3072$ correspondingly.
 The data exhibits a clear linear relationship, allowing us to perform a regression analysis and model the results using a linear function.
-By employing the method of least squares approximation, we obtain the subsequent functions as Equation 5 and 6:
+By employing the method of least squares approximation, we obtain the subsequent functions as Equation 6 and 7:
 $\mathcal{G}_{halving,2048}(\tau)=160604.41\cdot\tau-157178.66$ (6)
@@
 5.1 Theoretical Analysis on Proof Generation for Gas Usage Optimization
-The optimized proof generation in this research means minimizing the verification gas cost in EVM. We can define the total gas cost function $g_{total}$ using the previous regression results.
-$\mathcal{G}_{total}(\tau,\delta)=\mathcal{G}_{calldata}+\mathcal{G}_{halving}+\mathcal{G}_{exp}+C$ (7)
-$=\big{(}\alpha\cdot(\tau-\delta)+c_{1}\big{)}+\big{(}\beta\cdot(\tau-\delta)+c_{2}\big{)}+\big{(}\gamma\cdot 2^{\delta}+c_{3}\big{)}+C$ (8)
+The optimized proof generation in this research means minimizing the verification gas cost in EVM.
+We can define the total gas cost function $g_{total}$ using the regression results (Eq.
+27) in the previous section.
+$\mathcal{G}_{\text{total}}(\tau,\delta)=\mathcal{G}_{\text{calldata}}+\mathcal{G}_{\text{halving}}+\mathcal{G}_{\text{exp}}+C$ (8)
+$=\big{(}\underbrace{\alpha\cdot(\tau-\delta)+c_{1}}_{\mathcal{G}_{\text{calldata}}}\big{)}+\big{(}\underbrace{\beta\cdot(\tau-\delta)+c_{2}}_{\mathcal{G}_{\text{halving}}}\big{)}+\big{(}\underbrace{\gamma\cdot 2^{\delta}+c_{3}}_{\mathcal{G}_{\text{exp}}}\big{)}+C$ (9)
 $=(\alpha+\beta)(\tau-\delta)+\gamma\cdot 2^{\delta}+C^{\prime}$ (10)
@@
 Proof.
-See Appendix B. ∎
+See C. ∎
 Also, we denote the minimizing $\delta$ as
@@
 From Theorem 1, we can get the unique $\delta_{m}$ optimizing the gas cost.
-Applying all the regression results in Equation 1, 2, 3, 4, 5, and 6, we get the optimizing $\delta_{m}\approx 9.43$ for $\lambda=2048$ and $\delta_{m}\approx 9.15$ for $\lambda=3072$ respectively.
+Applying all the regression results in Equation 2, 3, 4, 5, 6, and 7, we get the optimizing $\delta_{m}\approx 9.43$ for $\lambda=2048$ and $\delta_{m}\approx 9.15$ for $\lambda=3072$ respectively.
 In practice, $\delta$ for implementation must be an integer.
@@
 Proof.
-See Appendix C. ∎
+See D. ∎
 Since $\log_{2}(\ln 2)\approx-0.52$, the optimal integer value of $\delta$ is 9 for both $\lambda=2048$ and $3072$, as the 1st case of Corollary 2.
 5.2 Optimized Verification Gas Cost
-Figures 4(a) and 4(b) depict the total gas cost for Pietrzak VDF verification at $\lambda=2048$ and $\lambda=3072$ respectively.
+Figures 3(a) and 3(b) depict the total gas cost for Pietrzak VDF verification at $\lambda=2048$ and $\lambda=3072$ respectively.
 Consistent with the theoretical predictions in Theorem 1, the gas usage graph shows a decline to a minimal point before increasing sharply.
@@
 25 4958138 9.63
-6 Conlusion and Future Works
+6 Discussion
+This section examines the broader implications and feasibility of implementing Pietrzak VDF verification on Ethereum.
+First, we consider the feasibility of the observed gas costs and discuss how frequent verification might remain economically viable on both Ethereum’s Layer 1 and Layer 2.
+Next, we address the lack of an established economic model for VDFs and the challenges it poses for incentivizing participants in decentralized environments.
+Lastly, we look beyond the EVM and outline how the findings could extend to more general computing contexts, suggesting avenues for future research and optimization.
+6.1 Practical Feasibility of Our Observed Costs.
+From Section 5, we note that even with an optimal parameter of $\delta=9$, the gas cost for verifying a Pietrzak VDF can reach a few million gas units.
+On Ethereum Layer 1, this is still somewhat expensive but may be acceptable for security-critical operations that are not executed frequently (e.g., infrequent random-beacon generation or governance steps).
+However, in Layer 2 environments with significantly reduced gas prices, these same verification procedures can become much more practical and cost-effective, especially for repeated or more frequent VDF verifications.
+By striking this balance—an optimal $\delta$ and a cheaper execution environment—our approach could feasibly support advanced use cases on EVM-compatible rollups, offering robust cryptographic guarantees at a fraction of Layer 1 costs.
+6.2 Lack of Economic VDF Protocols in Ethereum
+Integrating VDFs into smart contracts offers significant potential to enhance blockchain security and reliability, yet faces major challenges due to the lack of a robust economic model tailored to their unique operational needs, such as incentivizing both provers and verifiers and managing commit-reveal-recover protocols.
+This absence hampers practical deployment, as it fails to compensate those who contribute significant computational resources, deterring potential contributors and affecting scalability and effectiveness in real-world applications.
+Additionally, the decentralized nature of blockchain necessitates economic strategies that support technical operations and ensure sustainability without central governance, promoting honest participation and deterring dishonesty to maintain integrity and fairness within the blockchain ecosystem.
+Thus, innovative economic models specifically designed for decentralized environments are crucial to make VDFs both technically effective and economically viable.
+6.3 Future Works on a General Environment
+Through this research, we investigated the efficient implementation of Pietrzak VDF verification within the Ethereum Virtual Machine (EVM) environment and observed results that align both theoretically and experimentally.
+During our study, we identified significant effects of generalized proof generation and verification, which were also included within the predictable realms of theoretical analysis.
+We anticipate that such an approach could be applicable not only within the EVM but also in more general computing environments.
+Although resource pricing in general systems would differ from the gas costs in the EVM, with elements like the repetition of halving protocols and resource consumption proportional to the proof length being predictable, the costs of exponentiation in VDF ($2^{2^{T}}$) are expected to be exponential.
+Therefore, based on the outcomes of this study, further research on optimizing Pietrzak VDFs in general computing environments appears reasonable.
+7 Conclusion and Future Works
 Our study demonstrates that Pietrzak’s VDF can be effectively optimized for the EVM, reducing gas costs from 4M to 2M and minimizing proof size to under 8 KB with a 2048-bit RSA key.
@@
 We plan to implement on-chain Wesolowski VDF verification both directly and indirectly.
-For indirect implementation, a hash-to-prime function could be executed in a Truebit-style proof-of-fraud system.
-For direct implementation, refining the hash-to-prime function to be viable and secure on the EVM is essential.
+For indirect implementation, we suggest executing the hash-to-prime function in an off-chain, challenge-based verification system (inspired by Truebit [^4] Truebit [30] is a protocol that outsources computation off-chain while allowing anyone to challenge an allegedly correct result within a specific dispute window.
+If the challenge proves the result incorrect, a fraud proof is submitted on-chain, ensuring only correct computations are ultimately accepted.).
+This approach enables a “proof-of-fraud” mechanism, where a claimed correct computation can be disputed during a challenge period, thereby minimizing on-chain costs without compromising correctness.
 Appendix A Code availability
@@
 You can access the repository at the following link: https://github.com/usgeeus/Pietrzak-VDF-solidity-verifier
-Appendix B Proof of Theorem 1
+Appendix B Expected Gas Cost for Hash-to-Prime in Smart Contracts
+In Wesolowski’s VDF, implementing the hash-to-prime verification on-chain requires an algorithm capable of determining whether numerous candidate values are prime.
+As efficient probabilistic methods for prime checking, we considered Miller–Rabin primality test with 11 iterations, and the Baillie–PSW primality test which uses a single iteration of Miller–Rabin to base 2 and a Lucas primality test [31].
+For evaluation, we tested 6,100 randomly selected primes over $2^{255}$.
+On average, verifying a single prime consumed 21,446 gas for Miller–Rabin test and 45,912 gas for Baillie–PSW test.
+All prime number test cases, test scripts, and results are available in the repository referenced in A.
+In the worst case, the number of prime candidates can approach $\bigl{(}\log x\bigr{)}^{2}$ for large $x$, following Shanks’s upper bound on gaps between consecutive primes [32].
+Since we are dealing with the maximum number $12^{256}-$ 1, we set $x=2^{256}$, giving $\bigl{(}\log(2^{256})\bigr{)}^{2}\approx 5939$.
+Verifying 5939 candidate primes, each requiring about 21,446 gas for Miller–Rabin test or 45,912 gas for Baillie–PSW test, would quickly exceed the gas limit and render a purely on-chain implementation impractical.
+Moreover, even a single Miller–Rabin check can cost about \$0.28 (at 5 gwei) to \$1.12 (at 20 gwei) with ETH at \$2600, so multiplying by 5939 quickly runs into thousands of dollars, further underscoring the infeasibility of on-chain hash-to-prime.
+Appendix C Proof of Theorem 1
 Proof.
@@
 Thus, $\mathcal{G}_{total}(\tau,\delta)$ indeed has one unique minimum at $\delta$ within the specified range of $\tau>\delta\geq 0$, completing the proof. ∎
-Appendix C Proof of Corollary 2
+Appendix D Proof of Corollary 2
 Proof.